### v1

# thanks: current affiliation: IBM Quantum

Abstract
Quantum error correction (QEC) codes are necessary to fault-tolerantly operate quantum computers. However, every such code is inherently limited by its inability to detect logical errors. Here, we propose and implement a method that leverages dynamical decoupling (DD) to drastically suppress logical errors. The key to achieving this is to use the logical operators of the QEC code as DD pulses, which we refer to as logical dynamical decoupling (LDD). The resulting hybrid QEC-LDD strategy is in principle capable of handling arbitrary weight errors. We test QEC-LDD using IBM transmon devices and the $[[4,2,2]]$ code, demonstrating performance that significantly exceeds the capabilities of using either this code or DD in isolation. We present a method that allows for the detection of logical errors affecting logically encoded Bell states, which, in this case, arise primarily from crosstalk among physical qubits. Building on this, we experimentally demonstrate high-fidelity entangled logical qubits.
Quantum error correction (QEC) [1, 2, 3, 4] is fundamental to the realization of fault-tolerant quantum computation [5, 6, 7, 8], ensuring the preservation of quantum information undergoing error processes during computation and storage [9, 10]. Numerous successful experimental demonstrations of QEC across various platforms have been reported over the years [11, 12, 13, 14, 15], with the scale and pace accelerating recently towards genuine fault tolerance [16, 17, 18, 19, 18, 20, 21, 22, 23, 24].
Fault-tolerant quantum computing will require the stability of logical qubits over the potentially hours-long timescales of quantum algorithms solving utility-scale problems [25, 26, 27]. Above threshold, over such timescales low-weight physical errors may transform into logical errors of weight higher than any deployed fixed-distance QEC code can handle. This can become a concern even below threshold since long-range spatial and temporal correlations may develop that violate the assumptions underlying fault tolerance theory [28, 29, 30, 31, 32].
Conventionally, suppressing higher-weight errors requires increasing the QEC code distance, e.g., by means of code concatenation [33, 34], or increasingly larger codes such as is done with surface codes [35, 36], color codes [37], or quantum low-density parity-check codes [38]. Although effective, these strategies result in significant overhead, substantially increasing the number of physical qubits required and the time required to decode and correct errors.
Here, we propose and demonstrate a low-cost method that combines a fixed-distance QEC or quantum error detection (QED) code with dynamical decoupling (DD) [39, 40, 41, 42]. This hybrid method, which we call QEC-LDD, can handle arbitrary-weight errors.
Here, LDD stands for ‘logical dynamical decoupling’; the decoupling sequence is implemented using the logical operators of the QEC code as pulses.
We prove a corresponding QEC-LDD theorem which, informally, posits that the most general set of errors acting on an $[[n,k,d]]$ code, namely errors with weights up to $n$, can all either be suppressed using LDD or corrected using the code, for any $d\geq 1$. This seemingly paradoxical result is enabled by virtue of the fact that the LDD sequence suppresses all logical errors as well as some physical errors, thus leaving a much smaller subset of errors for the code to handle, which turn out to all be correctable. Notably, these are not the errors the code was originally designed to correct; indeed, the result holds in particular also for purely error detecting codes. The QEC-LDD theorem is a formal result, in the sense that QEC-LDD, if implemented literally as described in the theorem, is not expected to outperform a pure DD protocol (absent any QEC) on $k$ physical qubits. Nevertheless, the QEC-LDD theorem serves as a useful framework for less ambitious protocols that do not attempt to remove all errors but instead focus on the more physically relevant subset. We discuss and illustrate this conclusion in detail in this work, in the context of a transmon-based demonstration of entangled logical qubits.
Standard, ‘physical’ DD, where DD pulses act not at the logical level but at the physical qubit level, has recently shown great progress, improving the fidelity of storing quantum states [43, 44, 45, 46, 47, 48, 49], circuits [50, 51, 52], and even the performance of entire algorithms [53, 54, 55, 56]. Furthermore, physical DD can be seamlessly combined with fault-tolerant QEC [57, 58] and several recent QEC experiments have used physical DD profitably [18, 21, 19, 59, 20, 60].
However, such pulse sequences can introduce additional errors due to control imperfections in the pulses and due to crosstalk, potentially overshadowing the benefits of DD. To address this challenge, we design our LDD sequences to be robust against such control errors [61, 62] and crosstalk [63, 64, 65, 66, 67], ensuring that the advantages of logical error suppression are not compromised. This robust design allows our LDD implementation to enhance the protection of the code space as intended, effectively targeting both logical and control errors. We demonstrate the practical utility and advantages of this approach using $127$-qubit IBM quantum processors.
The workhorse in our demonstrations is the $[[4,2,2]]$ quantum error detection code, whose two logical qubits we use to prepare logically encoded Bell states. To demonstrate the effectiveness of LDD in suppressing logical errors, we first need a method to detect such errors. However, this code is constrained by its low distance and is incapable of detecting prevalent logical errors, including $ZZ$ crosstalk errors. To overcome this limitation, we design experiments in which, through the use of logically encoded Bell states, we nevertheless unequivocally detect the occurrence of logical errors for a known input logical state. This is then followed by the implementation of various versions of LDD, each employing different subsets of logical operators to construct the sequence. These implementations effectively demonstrate significant suppression of logical errors and substantial improvement in the fidelity of the code space.
The structure of this paper is as follows. In Section I we provide pertinent background on the error model and DD from the perspective of group averaging. We discuss standard Pauli-group DD as well as DD implemented using the stabilizers and or logical operators of a stabilizer code. In Section II we state and prove the QEC-LDD Theorem. We also illustrate the theorem using the example of the $[[4,2,2]]$ code. We then turn to our experimental demonstration of high-fidelity logical Bell states in Section III. This section describes our experimental design and methodology, evidence that both physical DD and error detection with postselection improve logical Bell state fidelity, and finally the evidence that the hybrid QEC-LDD strategy is significantly better.
We conclude in Section IV and provide supporting technical details in the Appendix.
I Background
I.1 Quantum codes and logical errors
An $[[n,k,d]]$ code encodes $k$ logical qubits into $n$ physical qubits with distance $d$ [68]. The weight of an error is the number of physical qubits it affects simultaneously. The highest-weight errors an $[[n,k,d]]$ code can correct or detect are $\lfloor\frac{d-1}{2}\rfloor$ and $d-1$, respectively. A $d=2$ code is a pure quantum error detection (QED) code, and a distance-$d$ QEC code can always be used as a QED code for errors of weight $<d$. The code distance is the minimum number of physical qubits that must experience an error to cause an undetectable logical error, i.e., an error forming a logical operation inside the code space. These logical errors can either be inherently present or result from the accumulation of lower-weight errors over time. For example, in the context of superconducting qubits with fixed-frequency couplers, a prevalent challenge is $ZZ$ crosstalk [69, 70], which inherently introduces weight-two errors that can present as logical errors for distance-$2$ codes.
I.2 Logical dynamical decoupling
Let $\mathcal{P}_{n}=\{\pm P_{j},\pm iP_{j}\}_{j=0}^{4^{n}-1}\subset\mathsf{U}(2^{n})$ be the full Pauli group with $4^{n+1}$ elements generated by $\{I,X,Y,Z\}^{\otimes n}$ and the phase factors {±$\{\pm 1,\pm i\}$. Let $\tilde{\mathcal{P}}_{n}=\mathcal{P}_{n}/\{\pm 1,\pm i\}\subset\mathsf{U}(2^{n})/\mathsf{U}(1)$ be the Abelian factor group with $4^{n}$ elements (here $\mathsf{U}(m)$ is the group of unitary operators on $\mathbb{C}^{m}$).
Consider the most general ‘total decoherence’ system-bath interaction for $n$ qubits:
$H_{SB}=\sum_{j=0}^{4^{n}-1}\alpha_{j}P_{j}\otimes B_{j}.$ (1)
We let $P_{0}=I^{\otimes n}$, the $n$-qubit identity operator. The pure-bath Hamiltonian $P_{0}\otimes H_{B}$ is harmless.
In its simplest form, accounting only for first-order decoupling and ignoring pulse errors, DD theory can be understood as group symmetrization [71, 72]. Choose a decoupling group $\tilde{\mathcal{G}}$ that consists of a set of unitary transformations (pulses) $g_{j}$ acting purely on the system: $\tilde{\mathcal{G}}=\left\{g_{j}\right\}_{j=0}^{\left|\tilde{\mathcal{G}}\right|-1}$, where $g_{0}=P_{0}$. We assume that each such pulse is a unitary operator defined up to a global phase, i.e. $\tilde{\mathcal{G}}\subset\mathsf{U}(2^{n})/\mathsf{U}(1)$; however, to simplify the notation we write, e.g., $g_{0}=P_{0}$ instead of $g_{0}=[P_{0}]$ (the equivalence class of $P_{0}$ up to a global phase). Similarly to $\mathcal{P}_{n}$ and $\tilde{\mathcal{P}}_{n}$, we assume that $\tilde{\mathcal{G}}$ can be written as $\tilde{\mathcal{G}}=\mathcal{G}/\{\pm 1,\pm i\}$ for some group $\mathcal{G}\subset\mathsf{U}(2^{n})$. In particular, we assume that the pulses are instantaneous and cycle over all elements of the group $\tilde{\mathcal{G}}$. This yields the DD pulse sequence:
$U(T)=\prod_{j=0}^{\left|\tilde{\mathcal{G}}\right|-1}g_{j}^{\dagger}f_{\tau}g_{j}=e^{-iT(\langle H_{SB}\rangle_{\mathcal{G}}+P_{0}\otimes H_{B})}+O(T^{2}).$ (2)
Here, $\tau$ is the pulse interval, $T=|\tilde{\mathcal{G}}|\tau$ is the total duration of one repetition of the sequence, $f_{\tau}=e^{-i\tau H_{SB}}$ is the free-evolution unitary, and $\langle H_{SB}\rangle_{\mathcal{G}}$ is the group-symmetrized system-bath Hamiltonian, where
$\langle A\rangle_{\mathcal{G}}\equiv\frac{1}{|\mathcal{G}|}\sum_{g\in\mathcal{G}}g^{\dagger}Ag=\frac{1}{|\tilde{\mathcal{G}}|}\sum_{g\in\tilde{\mathcal{G}}}g^{\dagger}Ag$ (3)
is the projection of the arbitrary bounded system-bath operator $A$ onto the subalgebra of operators that commute with every element of $\mathcal{G}$ (i.e., its commutant). We say that $\tilde{\mathcal{G}}$ decouples $H_{SB}$ (to first order) if $\langle H_{SB}\rangle_{\mathcal{G}}=cI$ where $c$ is a constant, including zero.
For example, $\tilde{\mathcal{G}}=\tilde{\mathcal{P}}_{n}$ decouples an arbitrary $n$-qubit system-bath Hamiltonian since then $\langle H_{SB}\rangle_{\mathcal{G}}=0$, albeit at a sequence time cost of $T=4^{n}\tau$ [40]. To illustrate this, consider $n=1$; the most general system-bath Hamiltonian of a single qubit is $H_{SB}=\sum_{i=0}^{3}\alpha_{i}P_{i}\otimes B_{i}$ with $P_{i}\in\{I,X,Y,Z\}$. If $\tilde{\mathcal{G}}=\{I,X,Y,Z\}$, then $U(T)$ simplifies into the well-known XY4 sequence $U(4\tau)=Yf_{\tau}Xf_{\tau}Yf_{\tau}Xf_{\tau}$ [73], and indeed, $\langle H_{SB}\rangle_{\mathcal{G}}=0$.
Higher-order decoupling sequences are known that achieve suppression up to arbitrary order $q$ [i.e., leaving an $O(T^{q+1})$ unsuppressed error term in $U(T)$[74, 75, 76, 77, 78], but for simplicity, in this work we restrict our attention to first-order sequences. Note that a common misconception is that DD is not effective against purely Markovian noise [79]; however this is not the case, essentially since even in the Markovian limit the bath can have a nonzero correlation time [80, 81, 82, 83].
[…]
For a code $\mathcal{C}\subset(\mathbb{C}^{2})^{\otimes n}$, the stabilizer group $\mathcal{S}$ is uniquely defined as
$\mathcal{S}=\{P\in\mathcal{P}_{n}:\forall\ket{\psi}\in\mathcal{C},\;P\ket{\psi}=\ket{\psi}\}.$ (4)
Its centralizer is also uniquely defined. However, the group of canonical logical operators is not unique: one can choose arbitrary stabilizers $S_{j}^{X},S_{j}^{Z}\in\mathcal{S}$ ($j=1,\dots,k$) and multiply the generators $\bar{X}_{j},\bar{Z}_{j}$ by those to obtain other choices of the group $\mathcal{L}$.
It is always possible to write the system-bath interaction [Eq. 1] as
$H_{SB}=H_{SB}^{\mathcal{S}}+H_{SB}^{\mathcal{L}}+H_{SB}^{\mathcal{E}},$ (5)
grouping together terms with similar action on the code space $\mathcal{C}$. Altogether, this expression has $4^{n}$ linearly independent terms corresponding to the elements of $\tilde{\mathcal{P}}_{n}$. Specifically, $H_{SB}^{\mathcal{S}}$ collects all the distinct $2^{n-k}$ Pauli operators $P_{i}\in\mathcal{S}$ (including $P_{0}=I$); such terms act as the identity operator on the code space and, therefore, are harmless. $H_{SB}^{\mathcal{L}}$ collects all $2^{n-k}(4^{k}-1)$ Pauli operators in $(\mathcal{S}\mathcal{L}\setminus\mathcal{S})/\{\pm 1,\pm i\}$; these terms leave the code space (and each of the other syndrome spaces) invariant but cause harmful logical errors, since by construction, the code cannot detect such errors. Finally, $H_{SB}^{\mathcal{E}}$ collects all remaining $2^{n+k}(2^{n-k}-1)$ terms in $\tilde{\mathcal{P}}_{n}\setminus(\mathcal{S}\mathcal{L}/\{\pm 1,\pm i\})$, which correspond to potentially detectable errors: given the value of the syndrome one can either detect all of them or correct some of them.
A key observation is that the decoupling group is arbitrary and may, in particular, be formed from the stabilizer group $\mathcal{S}$ and/or the group of canonical logical operators $\mathcal{L}$ of a stabilizer code $\mathcal{C}$, or even a subset of either one [87, 72, 88, 89, 58, 90]. One approach for selectively combining DD with QEC, which leverages this flexibility, is to use DD only to suppress the errors between or within the $2^{n-k}$ syndrome spaces, including the code space $\mathcal{C}$ (i.e., errors in $(\mathcal{S}\mathcal{L}\setminus\mathcal{S})/\{\pm 1,\pm i\}$). Errors causing transitions between the syndrome spaces (i.e., elements of $\tilde{\mathcal{P}}_{n}\setminus(\mathcal{S}\mathcal{L}/\{\pm 1,\pm i\})$) are the ones the code aims to detect or correct. The former are decoupled by $\mathcal{G}=\mathcal{L}$ while the latter are decoupled by $\mathcal{G}=\mathcal{S}$. For any choice of $\mathcal{G}$, the decoupling condition is that there exists at least one element of $\mathcal{G}$ that anticommutes with a given error. Based on these properties, it was shown in Ref. [58] that combining the full set of $n+k$ stabilizer and canonical logical operator generators to generate the DD group $\mathcal{G}=\mathcal{S}\times\mathcal{L}$, optimally achieves the goal of decoupling the errors between or within the $2^{n-k}$ syndrome spaces, in the sense of minimizing the length of the DD pulse sequence. This is known as ‘stabilizer-logical DD’ (SLDD) and incurs a single sequence cost of $T=2^{n+k}\tau$, instead of $4^{n}\tau$ in the case of full Pauli group decoupling. The savings are substantial when $k\ll n$. Adding QEC syndrome measurements and recovery operations based on syndrome measurements then detects the remaining errors between syndrome spaces that DD did not suppress [i.e., the $O(T^{q+1})$ terms in the case of a $q$’th-order DD sequence] or successfully corrects some of them, resulting in a hybrid QEC-DD strategy.
Figure 1: (a) Schematic of the experimental design: we encode a logical Bell state $\chi$ using the $[[4,2,2]]$ code and let it undergo either free evolution or LDD for a duration $\tau$. We then run the unencoding circuit for $\chi^{\prime}$ and measure the resulting bitstring. The encoding and unencoding circuits together take up to $5\mu$s. (b) The logical error set that relates the four Bell states as a result of a logical operator acting on one ($i=1$ or $2$) of the two logical qubits. (c) Interpretation of different bitstrings for an encoded state ($U_{\rm{enc},\chi}$), depending on the unencoding ($U_{\rm{enc},\chi^{\prime}}^{\dagger}$) circuit used. The color code indicates that the encoded state has undergone either no error (black) or one of the logical errors (red, blue, green) indicated in (b). The DD sequences used in our experiments are (d) Robust Logical XX (RLXX), (e) Robust Logical XY4 (RLXY4), and (f) physically Staggered XY4 (SXY4). Each panel shows both the physical implementation (top) and the logical interpretation of each sequence. A tilde in (d) and (e) denotes a negative $\pi$ rotation, i.e., $X=R_{X}(\pi)=\exp(-i\pi\sigma^{x})$ $\tilde{X}=R_{X}(-\pi)=\exp(i\pi\sigma^{x})$, and likewise for $Y$.
II The QEC-LDD Theorem
The main weakness of SLDD is that it still incurs a cost that is exponential in the number of physical qubits $n$. Here, we present a different hybrid approach, shifting more of the burden for handling correctable errors to the QEC code and using only the logical operators as DD pulses $g_{j}$. We call this approach ‘logical DD’ (LDD). Compared to SLDD, this approach has the desired effect of reducing the single sequence time-cost to $T=4^{k}\tau$ for first-order decoupling, independent of $n$. Table 1 summarizes the relative resource cost of full Pauli group DD, SLDD, and LDD.
Comparative DD resource cost
Decoupling group $\mathcal{G}$ of pulses suppresses all errors afflicting code space / requires QEC
Pauli group $\tilde{\mathcal{P}}_{n}$ $2^{2n}$ Yes / No
SLDD $2^{n+k}$ Yes / No
LDD $2^{2k}$ No / Yes
Table 1: The comparative resource cost for various DD methods in terms of the number of pulses required to suppress the most general system-bath Hamiltonian $H_{SB}$ [Eq. 5] acting on $n$ qubits, and whether active error correction is required in addition. Full Pauli group DD and SLDD suppress all error mechanisms which can decohere code states, while LDD does not, as it leaves dealing with these error mechanisms for the QEC cycle. Consequently, LDD does not scale with $n$ but with $k$, whereas Pauli group DD and SLDD both scale with $n$.
II.1 Theorem statement and proof
The LDD sequence is purely responsible for decoupling the logical errors acting within each syndrome space, but unlike SLDD, it no longer decouples some of the physical errors between the syndrome spaces. Correcting these errors is accomplished by syndrome measurements in conjunction with error recovery.
For each code $\mathcal{C}$ there is a choice of which errors to correct: for each of the $12^{n-k}-$ 1 nontrivial syndromes, one may choose a recovery operation; this choice determines which errors are successfully corrected and which result in a logical error. Once such a choice is made, we call the errors that are successfully corrected “correctable”.
We summarize this in the following theorem:
Theorem 1 (QEC-LDD).
Consider the most general system-bath Hamiltonian acting on $n$ qubits [Eq. 5]. Let $\mathcal{C}$ be an $[[n,k,d]]$ stabilizer code (with any $d\geq 1$). Let $\mathcal{L}=\langle i,\overline{X}_{j},\overline{Z}_{j}\rangle_{j=1}^{k}$ denote its group of canonical logical operators. Choose the decoupling group $\tilde{\mathcal{G}}=\mathcal{L}/\{\pm 1,\pm i\}$, henceforth the LDD group. Then $\tilde{\mathcal{G}}$ decouples all logical errors, that is, every term in $H_{SB}^{\mathcal{L}}$, as well as some errors in $H_{SB}^{\mathcal{E}}$. There is a choice of recovery operation such that all remaining non-decoupled errors in $H_{SB}^{\mathcal{E}}$ are correctable by $\mathcal{C}$. Conversely, for any recovery operation chosen in this way, only those errors that are not decoupled by $\tilde{\mathcal{G}}$ are correctable.
Proof.
We first establish when an error $E\in\mathcal{P}_{n}$ is decoupled. Let $\mathcal{H}=(\mathbb{C}^{2})^{\otimes n}$ be the Hilbert space of $n$ qubits. Then $\mathcal{C}\subset\mathcal{H}$ and all elements of $\mathcal{P}_{n}$ (and, hence, of $\mathcal{L}\subset\mathcal{P}_{n}$ and $\mathcal{S}\subset\mathcal{P}_{n}$) are unitary operators on $\mathcal{H}$. We let $\mathcal{G}=\mathcal{L}=\{\pm g_{j},\pm ig_{j}\}_{j=0}^{4^{k}-1}$. Consider a term in $H_{SB}$ corresponding to some $E\in\mathcal{P}_{n}$. It is decoupled if and only if there exists $g\in\mathcal{G}$ anticommuting with $E$. Indeed, then $g^{\dagger}Eg=-E$ and
$\langle E\rangle_{\mathcal{G}}=\frac{1}{\left|\mathcal{G}\right|}\sum_{h\in\mathcal{G}}h^{\dagger}Eh=\frac{1}{\left|\mathcal{G}\right|}\sum_{h\in\mathcal{G}}(gh)^{\dagger}E{gh}=-\langle E\rangle_{\mathcal{G}},$ (6)
so $\langle E\rangle_{\mathcal{G}}=0$. If there is no such $g$ then all elements of $\mathcal{G}$ commute with $E$ (because every two elements of $\mathcal{P}_{n}$ either commute or anticommute). Then all $g^{\dagger}Eg=E$ for all $g\in\mathcal{G}$, hence $\langle E\rangle_{\mathcal{G}}=E$ (i.e., the term is undecoupled). It follows that $\tilde{\mathcal{G}}$ decouples all logical errors, since every such error anticommutes with an element of $\mathcal{G}$. The errors in $H_{SB}^{\mathcal{E}}$ that anticommute with an element of $\mathcal{G}$ are also decoupled. All terms in $H_{SB}^{\mathcal{S}}$ are undecoupled since they all commute with $\mathcal{G}$.
The key observation in the proof is that the statement of the theorem is invariant under Clifford maps, which we now show.
Now consider any Clifford unitary $V\colon\mathcal{H}\to\mathcal{H}$. Let $\mathcal{C}^{\prime}=V\mathcal{C}$, $d^{\prime}\geq 1$ be the distance of the code $\mathcal{C}^{\prime}$, $\mathcal{G}^{\prime}=\mathcal{L}^{\prime}=V\mathcal{L}V^{\dagger}$, $\tilde{\mathcal{G}}^{\prime}=\mathcal{G}^{\prime}/\{\pm 1,\pm i\}$, $H_{SB}^{\prime}=VH_{SB}V^{\dagger}$, etc. For a recovery map $R_{s}$ (for every value $s$ of the syndrome) for $\mathcal{C}$ we can construct the corresponding recovery map $R_{s}^{\prime}=VR_{s}V^{\dagger}$ for $\mathcal{C}^{\prime}$. If the statement of the theorem is true for the original code $\mathcal{C}$, then it remains true for $\mathcal{C}^{\prime}$ because it only depended on the commutation relations between elements of $\mathcal{P}_{n}$ and those are preserved under the group homomorphism $P\mapsto VPV^{\dagger}$. By applying this observation to $V^{\dagger}$ instead of $V$ we observe that the converse also holds.
In 1 in Appendix C, extending earlier methods for stabilizer code encoding [91, 92, 93], we show that for any code $\mathcal{C}$ and any choice of the canonical logical operators $\mathcal{L}$ there is a Clifford unitary $V$ (which we may call “unencoding circuit”) such that the corresponding code $\mathcal{C}^{\prime}$ consists of states of the form |$\ket{\psi}\otimes\ket{0}^{n-k}$, where |$\ket{\psi}\in(\mathbb{C}^{2})^{\otimes k}$, and $\mathcal{L}^{\prime}=V\mathcal{L}V^{\dagger}\simeq\mathcal{P}_{k}$ consists of Paulis $P\in\mathcal{P}_{n}$ acting trivially on the last $n-k$ qubits. We call such code $\mathcal{C}^{\prime}$ a trivial code: its logical qubits can be chosen to coincide with the first $k$ physical qubits and stabilizer generators can be chosen to be $Z_{k+1},\dots,Z_{n}$ enforcing that the last $n-k$ qubits are set to 00, and $d^{\prime}=1$.
It follows that it is sufficient to prove the statement of the theorem for this trivial code. Hence, from now on we assume that $\mathcal{C}$ is the trivial code described above. $E\in\mathcal{P}_{n}$ is undecoupled if and only if it commutes with all Paulis acting on the first $k$ qubits. This is true if and only if $E$ acts trivially on the first $k$ qubits. (Equivalently, $E\in\mathcal{P}_{n}$ is decoupled if and only if acts non-trivially on one or more of the first $k$ qubits.) Measuring the stabilizers $Z_{k+1},\dots,Z_{n}$ is equivalent to measuring the last $n-k$ qubits. Any such measurement outcome $s=(s_{1},\dots,s_{n-k})$ corresponds to errors $E=SE_{s}$, where $E_{s}=\prod_{j=1}^{n-k}X_{k+j}^{s_{j}}$ and $S$ is an arbitrary stabilizer. A (unitary) recovery operation $R_{s}$ corrects such an error $E$ if and only if $R_{s}=S^{\prime}E_{s}$. In particular, if we apply $E_{s}$ corresponding to the measured syndrome $s$, we correct all non-decoupled errors, proving the first statement of the theorem. On the other hand, all such recovery maps act trivially on the first $k$ qubits, so if $E$ acts non-trivially on the first $k$ qubits (i.e. if $E$ is decoupled), it remains uncorrected by any such procedure. ∎
We emphasize a few key points:
  • The claim holds even for $d=2$ (or $d=1$), i.e., an error detecting code becomes an error correcting code subject to LDD.
  • In fact, the notion of correctable errors used here ignores the code distance altogether. For any distance, the QEC-LDD protocol has at least $3k$ non-correctable single-qubit errors (these are decoupled by LDD). [^1] Consider the group $\mathcal{L}$ and divide the physical qubits into three classes according to the action of the elements of $\mathcal{L}$ on a particular qubit. 1. Every element of $\mathcal{L}$ acts as $I$ on that qubit. 2. Every element of $\mathcal{L}$ acts as $I$ or one other Pauli ($X$, $Y$, or $Z$) on that qubit. 3. For every Pauli in $\{I,X,Y,Z\}$ there is $P\in\mathcal{L}$ which acts like that Pauli on that qubit. Let $l_{1},l_{2},l_{3}$ be the number of physical qubits in each class. Restrict every element of $\mathcal{L}$ to the qubits from class 3. Such a restriction does not change the commutation relations. Let the resulting group be $\mathcal{L}^{\prime}$. We have $\mathcal{P}_{k}\simeq\mathcal{L}\simeq\mathcal{L}^{\prime}$, which in turn is embedded in $\mathcal{P}_{l_{3}}$. Therefore, $l_{3}\geq k$. The number of uncorrectable single-qubit errors is $2l_{2}+3l_{3}\geq l_{3}\geq 3k$. E.g., for the $[[4,2,2]]$ code $l_{2}=0,l_{3}=3,k=2$, $2l_{2}+3l_{3}=9>6=3k$. For example, in the standard distance-$d$ surface code there is a weight $1$ error, which can be “corrected” to become a weight $12d-$ 1 error. That does not contradict the theorem, because that error is suppressed by LDD.
  • For each syndrome, the corresponding recovery operation is unique up to multiplication by stabilizers.
  • All decoupled errors are not correctable; it is the role of LDD to suppress such errors.
II.2 Illustration of the QEC-LDD Theorem: the $[[4,2,2]]$ code
To illustrate the QEC-LDD construction, we analyze the example of the $[[4,2,2]]$ code (see Appendix A for a review of this code). The stabilizer group of this code is $\mathcal{S}=\langle XXXX,ZZZZ\rangle$. A set of logical operators for the code can be chosen such that $\overline{X}_{1}=XIIX$, $\overline{X}_{2}=IIXX$, $\overline{Z}_{1}=IIZZ$ and $\overline{Z}_{2}=ZIIZ$, i.e., $\mathcal{L}=\langle i,XIIX,IIXX,IIZZ,ZIIZ\rangle$. Up to a global phase, $\{XIIX,IIXX,IIZZ,ZIIZ\}$ is the generator set of the $16$-element decoupling group $\tilde{\mathcal{G}}=\{g_{j}\}$. This LDD group suppresses every logical error $E$ in the $4\times 16=64$-element $\mathcal{S}\mathcal{L}/\{\pm 1,\pm i\}$: the group-averaged logical system-bath Hamiltonian $\langle H_{SB}^{\mathcal{L}}\rangle_{\mathcal{G}}$ vanishes, as every one of its terms is a logical operator that anticommutes with an element of $\mathcal{G}$.
The errors that are not suppressed are easily seen to be the $16$-element set =$\mathcal{E}=\{P\otimes Q\otimes P\otimes P:P,Q\in\{I,X,Y,Z\}\}$, since they commute with all elements of $\mathcal{G}$. Each such error is detectable because it anticommutes with at least one element of $\mathcal{S}$. Moreover, every such error is correctable since it is equal to the product of the stabilizer $P\otimes P\otimes P\otimes P$ and $I\otimes R\otimes I\otimes I$, where $R=Q^{\dagger}P$ can be determined (up to the global phase) from the value of the syndrome. The recovery operation is then the weight-one operator $I\otimes R\otimes I\otimes I$.
Ignoring phases, the Pauli group on $4$ physical qubits, $\tilde{\mathcal{P}}_{4}$, has $2564^{4}=$ 256 elements. Of these, $64$ are the elements of $\mathcal{S}\mathcal{L}/\{\pm 1,\pm i\}$, $16$ are the unsuppressed errors $\mathcal{E}$, which leaves $176$ other elements $[P]=\{\pm P,\pm iP\}\in\tilde{\mathcal{P}}_{4}$ such that corresponding $P\in\mathcal{P}_{4}$ anticommutes with at least one element of $\mathcal{G}$ and is, therefore, suppressed by it.
Appendix C illustrates the unencoding of the $[[4,2,2]]$ code. In this case, $\mathcal{P}_{2}$ decouples all errors on the first two (data) qubits and only $X_{3}$ or $X_{4}$ are left as undecoupled bit-flip errors on the last two qubits.
II.3 Avoiding the use of virtual-$Z$ gates
One of the limitations of our experimental setup is that physical $Z$ gates are unavailable: fixed-frequency transmons utilize virtual-$Z$ gates [94, 95]. Such gates are problematic for DD [95]. We now demonstrate that there are DD sequences satisfying the conditions of the QEC-LDD theorem which do not use any physical $Z$ pulses. To construct such a sequence, we first choose generators of $\mathcal{L}$ that do not involve physical $Z$ gates. An example choice would be $\mathcal{L}=\langle i,XIIX,IIXX,IIYY,YIIY\rangle$. One could follow the construction below with this choice. However, in order to improve the ability of the DD sequence to suppress $ZZ$ crosstalk we suggest a different choice. First, multiply the original generators of $\mathcal{L}$ (that is, $XIIX,IIXX,IIZZ,ZIIZ$) by the stabilizers $XXXX$, $XXXX$, $XXXX$, and $ZZZZ$, respectively, to obtain $\mathcal{L}^{\prime}=\langle i,IXXI,XXII,XXYY,IZZI\rangle$. Then, after some group operations, we can write the same group using a different set of generators containing no $Z$’s: $\mathcal{L}^{\prime}=\langle i,XXII,IIYY,IYIY,XIXI\rangle$. The DD sequence consisting of two repetitions of $XIXI$, $IYIY$, $XIXI$, $IIYY$, $XIXI$, $IYIY$, $XIXI$, $XXII$ then corresponds to the decoupling group /{±$\mathcal{L}^{\prime}/\{\pm 1,\pm i\}$, and involves no $Z$ gates.
This sequence was obtained as follows: denote the generators of $\mathcal{L}$ (or $\mathcal{L}^{\prime}$) as $h_{1},h_{2},h_{3},h_{4}$ and choose any Gray code — a sequence of bitstrings $a_{0},\dots,a_{15}$ such that neighboring bitstrings differ in only a single digit. Finally, pick $g_{j}=h^{a_{j}}$ (for $j=0,\dots,15$), where $a_{j}$ is interpreted as a multiindex (i.e. $g_{j}=\prod_{k=1}^{4}h_{k}^{(a_{j})_{k}}$). The pulses of the DD sequence are $g_{j}g_{j+1}$ for $j=0,\dots,15$.
However, in our experiments we found that a decoupling group generated by a subgroup of $\mathcal{L}$ and a subgroup of $\mathcal{S}$ works better in practice because of its robustness and ability to suppress crosstalk. The next section is dedicated to demonstrating experimentally how the hybrid QEC-LDD protocol results in a significant fidelity enhancement of entangled logical qubits.
No DD XY4 UR6 UR8 UR10 UR18 RGA8a
|$\ket{\Phi_{+}}$ $65.44\pm 0.24$ $83.94\pm 0.18$ $83.08\pm 0.19$ $83.11\pm 0.19$ $82.05\pm 0.19$ $81.43\pm 0.20$ $\boxed{84.20\pm 0.18}$
|$\ket{\Phi_{-}}$ $61.23\pm 0.24$ $\boxed{83.78\pm 0.18}$ $81.92\pm 0.19$ $82.63\pm 0.19$ $81.07\pm 0.19$ $80.81\pm 0.18$ $83.34\pm 0.18$
|$\ket{\Psi_{+}}$ $66.18\pm 0.23$ $\boxed{84.68\pm 0.18}$ $84.17\pm 0.18$ $83.37\pm 0.19$ $82.46\pm 0.19$ $81.21\pm 0.19$ $84.65\pm 0.18$
|$\ket{\Psi_{-}}$ $65.43\pm 0.23$ $\boxed{86.57\pm 0.17}$ $85.20\pm 0.18$ $85.53\pm 0.18$ $84.17\pm 0.18$ $83.44\pm 0.17$ $86.38\pm 0.17$
[…]
III Experimental design and methodology
The data for the experiments we report here were collected on two separate occasions from a total of $24$ sets of four-qubit experiments run on the ibm_kyiv [96] quantum processor. Dataset 1, using $14$ sets of qubits, was collected during the week of August 12th (2024). Dataset 2, using $10$ sets of qubits, was collected on March 1st and 2nd (2025). These $10$ four-qubit sets differ from those in dataset 1 in order to test the robustness of our results. The figure captions below specify the dataset. For each data point in our results, fidelity is independently calculated for each qubit set and then bootstrapped by resampling. The mean fidelity and standard deviation are derived from the bootstrapped data, where larger error bars indicate greater variability among the qubit sets.
The ibm_kyiv processor consists of coupled, fixed-frequency transmons [97]. Such qubits exhibit an always-on interaction between adjacent pairs [98]. This crosstalk gives rise to weight-2 error terms that correspond to logical errors with our $[[4,2,2]]$ code choice. Other errors of weight $\geq 2$, if present, would likewise correspond to logical errors. To demonstrate that LDD can suppress all logical errors, we design an experiment that enables the unequivocal detection of such errors. To this end, we use the two logical qubits of the $[[4,2,2]]$ code to create logical Bell states |$\ket{\Phi_{\pm}}=(\ket{\overline{00}}\pm\ket{\overline{11}})/\sqrt{2}$ and |$\ket{\Psi_{\pm}}=(\ket{\overline{01}}\pm\ket{\overline{10}})/\sqrt{2}$. The corresponding encoding circuits $U_{\text{enc},\chi}$, where $\chi\in\{\ket{\Phi_{\pm}},\ket{\Psi_{\pm}}\}$, create two copies of physical Bell states, followed by the application of a logical controlled-NOT ($\overline{\text{CNOT}}$) to generate the logical Bell states (see Appendix A).
To estimate the fidelity of a logical Bell state $\chi$, we start from the physical ground state |$\ket{0000}$, encode into the state of interest by applying $U_{\text{enc},\chi}$, and then let it evolve for some time $\tau$, either freely or subject to DD.
If after time $\tau$ we unencode by applying $U_{\text{enc},\chi}^{\dagger}$, then measuring the bitstring $0000$ would indicate that no error occurred.
If, instead, we obtain any of the odd-Hamming weight bitstrings, this would signal detection of a physical error, as such states are not in the codespace. The final possibility is that we obtain one of the other even-Hamming weight bitstrings $\{0110,1001,1111\}$, corresponding to a logical error. This procedure can be used to estimate the probability of finding the ground state as the empirical fraction of ground state measurement outcomes, which is also the fidelity of the logical Bell state $\chi$.
Figure 2: Circuit schematic for encoding |$\ket{\Phi_{+}}$ and unencoding |$\ket{\Phi_{-}}$ with physical and logical DD. Physical XY4 dynamical decoupling sequences (yellow boxes) are inserted into the idle gaps of the circuit. The code subspace is protected with LDD sequences. This is the DD-protected version of the circuit shown in Fig. 1(a).
Figure 3: Performance of LDD (dataset 1). The experiments encode the logical Bell states |$\ket{\Phi_{+}}$ in (a) and (b), and |$\ket{\Psi_{+}}$ in (c) and (d). We then unencode in the logical Bell state related through $\overline{Z}_{i}$ on either of the logical qubits; i.e., |$\ket{\Phi_{+}}\mapsto\ket{\Phi_{-}}$ in (a) and (b), and |$\ket{\Psi_{+}}\mapsto\ket{\Psi_{-}}$ in (c) and (d). In this setting, the frequency of $0110$ outcomes [(a) and (c)] is a measure of the fidelity of the originally encoded state, while $0000$ corresponds to the detection of a $\overline{Z}_{i}$ error [(b) and (d)]. Without any DD, the fidelity is low (green in (a) and (c)) and the probability of logical errors is high [green in (b) and (d)]. Standard LDD sequences (LXX and LXY4; dashed light red and blue) improve the fidelity and logical error probability at short times ($\lesssim 15\mu$s) but their performance declines at longer times: they exhibit coherent errors as seen in the oscillations of the corresponding curves. Robust sequences (RLXX and RLXY4; solid dark red and blue) are the top performers; their ability to suppress the $\overline{Z}_{i}$ error is particularly noteworthy. Error bars are $1\sigma$ standard deviation after bootstrapping the data.
To obtain the fidelity of a logical error, we first observe that each of the logical Bell states is related to the other three logical Bell states through a specific logical error: $\chi^{\prime}=\overline{O}\chi$, where $\overline{O}\in\{I,\overline{X}_{i},\overline{Y}_{i},\overline{Z}_{i}\}$ and $i=1$ or $2$; i.e., a logical operator that acts on either of the two logical qubits. We thus proceed as follows: instead of applying $U_{\rm{enc},\chi}^{\dagger}$, at $t=\tau$ we deliberately unencode into a different logical Bell state $\chi^{\prime}$ [see Fig. 1(a)]. In this manner, each of the even-weight bitstrings gives us a measure of the occurrence of one of the logical errors $\mathcal{E}_{\overline{O},\tau}$, where $\overline{O}$ is determined by the unencoding we choose [see Fig. 1(b,c)]. Using this methodology, we can detect the occurrence of different types of logical errors and quantify the associated error probability.
In more detail, each row in Fig. 1(c) shows a different initially encoded logical Bell state. Each column corresponds to one of the unencoding circuits. In each unencoding scenario, the different bitstrings indicate that the state being unencoded is either the initial logical Bell state or some other logical Bell state. For example, consider the case where we originally encode |$\ket{\Phi_{+}}$ and then unencode into |$\ket{\Phi_{-}}$ at $t=\tau$. If the state being unencoded is indeed |$\ket{\Phi_{-}}=\overline{Z}_{i}\ket{\Phi_{+}}$, then the $0000$ bitstring signals that a $Z$ logical error ($\overline{Z}_{i}$) has occurred. However, if no error has occurred, then due to unencoding into |$\ket{\Phi_{-}}$, the result should be the $0110$ bitstring. Generalizing, it is possible to detect the different logical errors using the experiments indicated in Fig. 1(c).
However, we specifically choose the unencoding so that it is always the $0000$ bitstring that corresponds to the occurrence of the logical error operator in which we are interested. We make this choice since |$\ket{0000}$ is the ground state of the system and therefore is robust against relaxation errors. This strongly increases the likelihood that the detected errors are purely logical and are unaffected by thermal relaxation. Note that due to this choice, the bitstring that corresponds to the fidelity $\mathcal{F}_{\chi}(t)$ of the prepared initial state varies [black color-coded bitstrings in Fig. 1(c)].
From here on, we use the notation $\chi\mapsto\chi^{\prime}$ to denote the procedure of preparing the encoded logical Bell state $\chi$ and unencoding into $\chi^{\prime}$, i.e., of using the encoding unitary $U_{\text{enc},\chi}$ and the unencoding unitary $U^{\dagger}_{\text{enc},\chi^{\prime}}$.
We note that logical state tomography is an alternative method for learning about the performance of LDD, and in particular, for certifying the entanglement of logical Bell states. However, the $[[4,2,2]]$ code imposes some additional challenges in performing logical tomography, which precluded its use in our work; see Appendix D for details.
III.1 Physical dynamical decoupling improves logical Bell state fidelity
We first show that we can substantially improve the logical Bell fidelity by padding the idle gaps of the encoding circuits with physical (as opposed to logical) DD sequences. An idle gap is a temporal circuit segment during which no gates are applied. Such gaps occur, e.g., when a pair of qubits is involved in a two-qubit gate that takes much longer than a single-qubit gate simultaneously being applied to another qubit; the latter is then idle after the completion of the single-qubit gate, while awaiting the completion of the two-qubit gate.
As shown in Fig. 2, we inserted various DD sequences into the idle gaps of Bell state circuits. This includes XY4 [40], universally robust sequences URn [62] for $n=6,8,10,18$, and RGA8a [61]; the former are illustrated along with robust variants in Fig. 1(d-f).
We performed the experiments (during the week of August 12th 2024) on fourteen different sets of qubits of ibm_kyiv and report the average fidelity values in Table 2. In these experiments, we encoded and unencoded the same state, i.e., used $U_{\rm{enc},\chi}$ along with $U_{\rm{enc},\chi}^{\dagger}$. We see that using DD, the encoding fidelity improves by $\sim 20\%$ for all four logical Bell states, with XY4 and RGA8a being the top performers (we attribute the lower performance of the longer URn sequences to pulse interference effects [95]).
In all, these results demonstrate that logical Bell state preservation can benefit significantly from physical DD. Having established the utility of physical DD, we proceed to combine it with LDD in the next section.
Figure 4: Detection probability (i.e., probability of the $0000$ bitstring) of logical $X$, $Y$, and $Z$-type errors (dataset 1). The lower probability of logical $X$ and $Y$-type errors signals that the main source of logical errors is $ZZ$ crosstalk. These experiments are performed without DD.
III.2 Logical error suppression and detection by LDD
For all our experimental results, ‘No DD’ refers to $[[4,2,2]]$ encoding without any physical or logical DD. In all other experiments, we use physical XY4 to pad all the encoding and unencoding circuits gaps, and combine them with various flavors of LDD. This choice allows us to clearly assess the improvements introduced by DD, compared to relying solely on the code’s error detection capabilities.
III.2.1 Logical $Z$ errors
We start by gradually increasing the time delay $\tau$ (up to 55 $\mu$s) between the encoding and unencoding without DD. This situation is relevant in the context of QEC experiments. For example, one could prepare an encoded qubit and then leave it idle while other logical operations are applied to other encoded qubits [19].
We perform the $\overline{Z}_{i}$ error detection using the logical Bell states |$\ket{\Psi_{+}}$ and |$\ket{\Phi_{+}}$ as discussed in Section III, and display the results in Fig. 3. As can be seen in Fig. 3(a) and (c), without LDD the free evolution fidelity (denoted as No DD) decays rapidly and exhibits $ZZ$ crosstalk-induced oscillations. Fig. 3(b) and (d) show that logical $Z$ errors accumulate over time. We next use two types of LDD sequences to suppress this effect: Logical XX (LXX) and Logical XY4 (LXY4). Physical-level schematics of these LDD sequences are shown in the boxed four-pulse sequences of Fig. 1(d) and (e), respectively (disregarding the tilde notation). Their logical-level counterparts are shown at the bottom of Fig. 1. We generate both LXX and LXY4 using the native logical operations of the $[[4,2,2]]$ code. For example, in the upper part of Fig. 1(d), reading the first column from top to bottom yields $XIXI=\overline{XX}$, which is the $\bar{X}\bar{X}$ column at the bottom of Fig. 1(d). Reading the pulse sequences from left to right, the staggering (appearance of delays as indicated by the identity operations) is deliberately introduced to suppress crosstalk at the physical level [65]; see Appendix B for more details.
The corresponding results are shown as dashed lines in Fig. 3. At short times ($\lesssim 15\mu$s), these results are better than those without DD because some errors (including $ZZ$ crosstalk) are suppressed by both LXX and LXY4. However, at longer times, the benefit is lost and, moreover, small oscillations appear that indicate the presence of coherent errors  [99]. To overcome this, we create robust versions of LXY4 and LXX by ensuring that all physical qubits in the code undergo physical DD sequences robust to pulse errors. Specifically, we use the universally robust (UR) sequence family [62], and ensure that each physical qubit undergoes a UR4 sequence, i.e., $X\tilde{X}\tilde{X}X$ or $Y\tilde{Y}\tilde{Y}Y$, where a tilde denotes an $X$ or $Y$ rotation by $-\pi$ instead of $\pi$. These robust versions, which we call RLXX and RLXY4, are the full sequences shown in Fig. 1(d) and (e). The performance of these sequences is shown by the solid lines in Fig. 3 and exhibits a significant improvement. Notably, the logical Bell state fidelities decay more slowly and without oscillations (panels (a) and (c)), and the logical $Z$ errors are strongly suppressed (panels (b) and (d)).
In addition, we apply the physical staggered XY4 (SXY4) sequence. This corresponds to applying a single XY4 sequence to each physical qubit but in a staggered manner to reduce crosstalk [65]. This sequence, shown in Fig. 1(f), also corresponds to an LDD sequence for the $[[4,2,2]]$ code (bottom of panel (f)). However, it is not a universal logical sequence, but rather a purely logical-$Z$-error suppressing sequence. Interestingly, Fig. 3 shows that SXY4 performs on par with RLXX and RLXY4. This finding signals that the dominant logical errors are of $Z$-type. We confirm this in the following.
[…]
III.2.2 Logical $X$ and $Y$ errors
So far, we have only discussed the detection and suppression of $Z$-type logical errors. In order to detect $X$ and $Y$-type logical errors, we proceed in analogy to the $\overline{Z}$ error detection procedure, but with a different unencoding step. Namely, we start by encoding the logical |$\ket{\Phi_{+}}$ state, and then unencode in either |$\ket{\Psi_{+}}$ to detect $X$-type errors or in |$\ket{\Psi_{-}}$ to detect $Y$-type errors. The results are shown in Fig. 4 where, for comparison, we have also included the no-DD |$\ket{\Phi_{+}}\mapsto\ket{\Phi_{-}}$ and |$\ket{\Psi_{+}}\mapsto\ket{\Psi_{-}}$ results shown in Fig. 3(b,d), which measure $Z$-type logical errors. It is clear from Fig. 4 that while unsuppressed logical errors $Z$ accumulate rapidly, logical errors of $X$ and $Y$-type grow much more slowly. This confirms that $ZZ$ crosstalk is the main source of logical errors.
III.2.3 LDD with postselection
Having implemented LDD, we can further improve the results by using the error detection capability of the $[[4,2,2]]$ code, which allows us to perform postselection. We do so by discarding any measurement outcome outside the logical basis (which would be the result of a physical error); i.e., we only keep measurement results corresponding to the bitstrings $\{0000,0110,1001,1111\}$.
[…]
III.2.4 Beyond breakeven and the state-of-the-art
Fig. 5 also includes results for physical (unencoded) Bell states. Here we show only the best Bell pair among all pairs we tested, both without DD and with staggered (crosstalk-robust [65]) XY4. The former (‘Best Physical No DD’) exhibits strong crosstalk-induced oscillations with an overall fidelity comparable to that of the mean logical encoded Bell pairs case without DD (No DD+PS). Adding SXY4 significantly improves the fidelity and outperforms even the mean fidelity of logical Bell pairs with LDD (RLXY4). This shows that LDD by itself is not better than working with physical qubits and a crosstalk-robust DD sequence. However, physical Bell pairs with SXY4 are significantly worse than RLXY4+PS, i.e., the case of LDD with postselection on the results of the $[[4,2,2]]$ code. This constitutes clear evidence of beyond-breakeven performance for our QEC-LDD strategy.
[…]
Figure 6: The percentage of data discarded per circuit as a function of the delay time between encoding and unencoding (dataset 2). While LDD suppresses logical errors, it also reduces certain physical errors, as evidenced by the lower data discard rate when using LDD (i.e., fewer physical errors are detected).
III.3 Physical error suppression and detection by LDD
The LDD sequences suppress not only logical errors but physical errors as well. To see this, consider, e.g., the LXX sequence [Fig. 1(d)]. As described in Section III.2.1, reading the pulse sequence vertically, each time step of this sequence operates in the logical subspace as $\overline{XX}$. Simultaneously, reading the sequence horizontally over a complete round of LDD, all four physical qubits undergo the physical XX sequence (i.e., $X-\tau-X-\tau$), which suppresses the set of physical errors $\{Y_{i},Z_{i}\}_{i=1}^{4}$. Similarly, for the RLXY4 sequence [Fig. 1(e)], each physical qubit undergoes UR4. which robustly [62] suppresses the sets of physical errors $\{Y_{i},Z_{i}\}_{i=1,3}$ (due to the $X\tilde{X}\tilde{X}X$ sequence) and $\{X_{i},Z_{i}\}_{i=2,4}$ (due to the $Y\tilde{Y}\tilde{Y}Y$ sequence).
One way to gauge the impact of this suppression of physical errors is shown in Fig. 6, which displays the percentage of discarded data per circuit as a result of postselection over $55\mu$s (out of $4000$ shots), comparing No DD to RLXY4 for the two logical Bell states |$\ket{\Phi_{+}}$ and |$\ket{\Psi_{+}}$. The results are obtained through bootstrapping across the ten sets of four qubits used in these experiments. Significantly less data are discarded with RLXY4 than without DD. Since discarded data correspond to the detection of physical errors, this means that LDD not only enhances the fidelity of logical Bell states by suppressing logical errors, but also reduces the occurrence of physical errors.
We can go further and use Algorithmic Error Tomography [54] to identify specific physical error types. For example, the $0100$ and $0010$ bitstrings correspond, respectively, to physical $Z_{2},Z_{4}$ and $Z_{1},Z_{3},$ errors in the |$\ket{\Phi_{+}}\mapsto\ket{\Phi_{-}}$ logical Bell state experiment. In Fig. 7 we show the corresponding relative bitstring counts, which are measures of detecting either of these errors. Additionally, we show the $0000$ and $0110$ relative bitstring counts, which correspond to the logical fidelity and the logical $Z$ error, respectively. Without LDD (dashed), we observe that the logical fidelity is low and both logical and physical errors increase and oscillate. With DD (solid), the RLXY4 sequence strongly suppresses both types of errors.
These results demonstrate that, as claimed, LDD sequences suppress both physical and logical errors.
Figure 7: The relative count of the $\{0000,0110,0100,0010\}$ bitstrings in the |$\ket{\Phi_{+}}\mapsto\ket{\Phi_{-}}$ experiments (dataset 2). The two even-weight bitstrings correspond to the logical fidelity ($0000$) and logical $Z$ error ($0110$), and the two odd-weight bitstrings correspond to the detection of physical $\{Z_{i}\}_{i=1}^{4}$ errors. Dashed (solid) lines correspond to no DD (RLXY4). The LDD sequence, in addition to suppressing the logical errors and improving the logical fidelity, also strongly suppresses the occurrence of these physical errors.
IV Discussion
The operation of QEC codes is adversely affected by the occurrence of logical errors that the code cannot detect or correct. Here, we have shown how to combine QEC with dynamical decoupling implemented in terms of the logical operators of the code, resulting in a hybrid QEC-LDD strategy that is significantly more effective than either QEC or LDD alone.
We designed our LDD sequences to simultaneously perform logical error suppression and to be robust DD sequences at the physical level, resistant to both control errors and crosstalk. Our results, using the $[[4,2,2]]$ code and IBM transmon qubits, demonstrate a beyond-breakeven fidelity of entangled logical qubits.
Our findings address a need along the path toward fault-tolerant quantum computation: keeping codes relatively small and nimble while still effectively handling logical errors.
Future research should aim to optimize LDD sequences tailored to specific codes and integrate QEC-LDD into quantum algorithms. Another interesting future direction is optimization of QEC-LDD for tunable-coupler transmon devices; we present preliminary results in Appendix E.
Acknowledgements.
This material is based upon work supported by, or in part by, the Intelligence Advanced Research Projects Activity (IARPA), under the Entangled Logical Qubits program through Cooperative Agreement Number W911NF23-2-0216, by the U.S. Army Research Laboratory and the U.S. Army Research Office under contract/grant number W911NF2310255, and by the Defense Advanced Research Projects Agency under Agreement HR00112230006. The views, opinions and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government. This research was conducted using IBM Quantum Systems provided through University of Southern California’s IBM Quantum Innovation Center. The views expressed are those of the authors and do not reflect the official policy or position of IBM or the IBM Quantum team.
Appendix A [[4,2,2]] code
The $[[4,2,2]]$ code is an error detection code that encodes $k=2$ logical qubits into $n=4$ physical qubits [102]. The stabilizer group is $\mathcal{S}=\{I,XXXX,YYYY,ZZZZ\}$. Defining the logical states as
[…]
a set of logical operators for the code can be defined such that $\overline{X}_{1}=XIIX$, and $\overline{X}_{2}=IIXX$, up to multiplication by a stabilizer element. Therefore, we have $\overline{XX}=XIXI$. Similar definitions apply to the logical $Z$ operators: $\overline{Z}_{1}=IIZZ$ and $\overline{Z}_{2}=ZIIZ$, allowing us to form the full logical Pauli group. Using the same definitions, we have $\overline{\text{CNOT}}_{12}=\text{SWAP}_{12}$ and similarly, $\overline{\text{CNOT}}_{21}=\text{SWAP}_{23}$, i.e., we can perform logical CNOTs by swapping the physical qubits.
Our particular choice of logical operators is motivated by the fact that $\overline{XX}$ and $\overline{YY}$ (i.e., the logical operators we use to implement the LDD sequences) have a natural staggering of their physical $X$ and $Y$ gates. In other words, in the implementation of the logical operators comprising LDD, nearest-neighbor qubits are always interleaved with an identity operation (e.g., $IXIX$ as opposed to $IXXI$). This is critical for nearest-neighbor crosstalk cancellation [65].
To encode the logical Bell states |$\ket{\Psi_{\pm}}$ and |$\ket{\Phi_{\pm}}$ using the two logical qubits of the $[[4,2,2]]$ code, consider:
|$\ket{\Phi_{\pm}}$ $=\frac{1}{\sqrt{2}}\big{(}\ket{\overline{00}}\pm\ket{\overline{11}}\big{)}$ (8a)
[…]
$=\overline{\text{CNOT}}_{21}\frac{1}{\sqrt{2}}\big{[}(\ket{00}\pm\ket{11})(\ket{00}\pm\ket{11})\big{]}.$ (8d)
Thus, we proceed by preparing two physical copies of the target Bell state on the four physical qubits, then apply $\overline{\text{CNOT}}_{21}$, which creates the intended Bell state. The |$\ket{\Psi_{\pm}}$ Bell states are prepared similarly.
Appendix B Creating robust logical sequences
As discussed above, we use the logical operators $\overline{XX}=XIXI$ and $\overline{YY}=IYIY\equiv YIYI$ (along with identity operators) to generate the LDD groups ${\mathcal{G}}_{{\text{LXY4}}}=\{I,\overline{XX},\overline{YY},\overline{ZZ}\}$, and ${\mathcal{G}}_{{\text{LXX}}}=\{I,\overline{XX}\}$. Cycling over the group elements creates the four-pulse sequences LXY4 and LXX shown in the red boxes of Fig. 1(d,e).
Since the main source of logical errors is $ZZ$ crosstalk, we optimize LDD so that it cancels such errors. Thus, in the case of LXX, we insert a stabilizer every other pulse-step to consistently apply a staggered sequence to all qubits. These sequences correspond to applying logical $X$ and $Y$ operators to both logical qubits.
To create the robust sequences, we mirror the original sequences (hence the eight-pulse sequences) and instead of $X$ and $Y$, we apply $\tilde{X}=R_{X}(-\pi)$ $\tilde{Y}=R_{Y}(-\pi)$, respectively, such that each sequence undergoes a robust pulse sequence at the physical level. We also apply staggered physical XY4 to each qubit as shown in Fig. 1(f). This sequence is also inherently robust at the physical level since each qubit receives an XY4 sequence, which is robust to pulse errors [62]. However, it is not a universal decoupling sequence at the logical level, as seen in the figure (i.e., it lacks the ability to decouple arbitrary single logical-qubit errors).
Appendix C The encoding/unencoding Lemma
Encoding of stabilizer codes is a familiar problem [91]. Here we present an explicit version of encoding circuits that appears to be simpler than other versions found in the literature [92, 93].
C.1 Lemma and proof
Lemma 1.
For any $[[n,k,d]]$ stabilizer code it is always possible to construct a unencoding circuit $V^{\dagger}$ such that $V^{\dagger}$ converts any group of canonical logical operators and a corresponding stabilizer group into single-qubit Pauli $X$ and $Z$ operators acting just on the first $k$ physical data qubits and into single-qubit Pauli $Z$ stabilizer operators on the remaining $n-k$ ancilla qubits.
Proof.
Our proof is constructive and relies on the basic Clifford group identities $\text{CNOT}_{i,j}(X_{i}X_{j})\text{CNOT}_{i,j}=X_{i}$ and $\text{CNOT}_{i,j}(Z_{i}Z_{j})\text{CNOT}_{i,j}=Z_{j}$, where $\text{CNOT}_{i,j}$ denotes the controlled-NOT gate with qubit $i$ as the control and qubit $j$ as the target. Note that this identity allows us to reduce the number of $X$ operators.
Consider first the case $k>0$.
  1. 1.
    Order the canonical logical operators and stabilizers as $\{\overline{X}_{1},\overline{Z}_{1},\dots,\overline{X}_{k},\overline{Z}_{k},S_{1},...,S_{n-k}\}$.
  2. 2.
    Ensure that $\overline{X}_{1}$ consists only of $X$ and $I$ physical operators: apply $H=(X+Z)/\sqrt{2}$ and $(X+Y)/\sqrt{2}$ to the qubits where $Z$ and $Y$, respectively, are present in the expression for $\overline{X}_{1}$.
  3. 3.
    Apply SWAP1,j if necessary to ensure $X_{1}$ is present in the expression for $\overline{X}_{1}$.
  4. 4.
    Apply $Z_{1}$ if necessary to cancel the sign if the sign was $-1$.
  5. 5.
    Apply CNOT1,j to eliminate all $X$ components of $\overline{X}_{1}$ except $X_{1}$.
  6. 6.
    Apply analogs of steps 2, 4, and 5 for $\overline{Z}_{1}$.
  7. 7.
    Now note that all the other operators contain $I_{1}$ (indeed, we have $\overline{X}_{1}=X_{1}$, $\overline{Z}_{1}=Z_{1}$, and all other operators commute with these two, hence they cannot contain $X_{1}$, $Z_{1}$, or $Y_{1}$), so the task is reduced to the case with $k:=k-1$ and $n:=n-1$; proceed recursively.
Next, consider the case $k=0$. We have stabilizer operators $S_{1},...,S_{n}$ on $n$ physical qubits and wish to transform them to $Z_{1},...,Z_{n}$.
  1. 1.
    Apply analogs of previous steps 2-5 to $S_{1}$ to transform it into $Z_{1}$.
  2. 2.
    Recursively call this procedure to transform the remaining $n-1$ components of the remaining $n-1$ stabilizers to $Z_{2}$, …, $Z_{n}$.
  3. 3.
    Apply CNOT1,j gates to cancel the $Z_{1}$ components of the remaining $n-1$ stabilizers (if they had one). Note that they could not have had $X_{1}$ or $Y_{1}$ components because otherwise they would not commute with $S_{1}$.
  4. 4.
    Apply $X_{j}$ to cancel the sign of the remaining stabilizers if necessary.
C.2 Illustration using the $[[4,2,2]]$ code
As an example, consider the $[[4,2,2]]$ code. The stabilizer is $\mathcal{S}=\langle X^{\otimes 4},Z^{\otimes 4}\rangle$, and we choose the canonical logical operators as $\mathcal{L}=\langle\overline{X}_{1}=X_{1}X_{4},\overline{X}_{2}=X_{1}X_{2},\overline{Z}_{1}=Z_{1}Z_{2},\overline{Z}_{2}=Z_{1}Z_{4}\rangle$. The corresponding code basis states are
|$\ket{\overline{00}}=\frac{1}{\sqrt{2}}\ket{0000}+\ket{1111}$ (9)
|$\ket{\overline{10}}=\frac{1}{\sqrt{2}}\ket{0110}+\ket{1001}$
|$\ket{\overline{01}}=\frac{1}{\sqrt{2}}\ket{0011}+\ket{1100}$
|$\ket{\overline{11}}=\frac{1}{\sqrt{2}}\ket{1010}+\ket{0101}.$
Following the steps in the proof of 1, we obtain the following series of transformations:
$\overline{X}_{1}$ $=X_{1}X_{4}\stackrel{{\scriptstyle\text{CNOT}_{14}}}{{\longmapsto}}X_{1}{\longmapsto}X_{1}{\longmapsto}X_{1}{\longmapsto}X_{1}{\longmapsto}X_{1}$
$\overline{Z}_{1}$ $=Z_{1}Z_{2}\longmapsto Z_{1}Z_{2}\stackrel{{\scriptstyle\text{CNOT}_{21}}}{{\longmapsto}}Z_{1}{\longmapsto}Z_{1}{\longmapsto}Z_{1}{\longmapsto}Z_{1}$
$\overline{X}_{2}$ $=X_{1}X_{2}{\longmapsto}X_{1}X_{2}X_{4}{\longmapsto}X_{2}X_{4}\stackrel{{\scriptstyle\text{CNOT}_{24}}}{{\longmapsto}}X_{2}{\longmapsto}X_{2}{\longmapsto}X_{2}$
$\overline{Z}_{2}$ $=Z_{1}Z_{4}\longmapsto Z_{4}{\longmapsto}Z_{4}{\longmapsto}Z_{2}Z_{4}\stackrel{{\scriptstyle\text{CNOT}_{42}}}{{\longmapsto}}Z_{2}{\longmapsto}Z_{2}$
$S_{1}$ $=X^{\otimes 4}{\longmapsto}X_{1}X_{2}X_{3}{\longmapsto}X_{2}X_{3}{\longmapsto}X_{2}X_{3}X_{4}{\longmapsto}X_{3}X_{4}$
$\quad\stackrel{{\scriptstyle H_{3}H_{4}}}{{\longmapsto}}Z_{3}Z_{4}\stackrel{{\scriptstyle\text{CNOT}_{34}}}{{\longmapsto}}Z_{3}$
$S_{2}$ $=Z^{\otimes 4}{\longmapsto}Z_{2}Z_{3}Z_{4}{\longmapsto}Z_{2}Z_{3}Z_{4}{\longmapsto}Z_{3}Z_{4}{\longmapsto}Z_{3}Z_{4}$
$\quad{\longmapsto}Z_{3}Z_{4}{\longmapsto}Z_{4}{\longmapsto}Z_{4}\ ,$ (10)
where we have only indicated the transformations once, for the operator that induces the transformation per the proof of 1; e.g., $\text{CNOT}_{14}$ is induced by $\overline{X}_{1}$ but is applied to every operator in the first timestep, etc.
Figure 8: Performance of LDD on for an encoded |$\ket{\Phi_{+}}$ Bell state on ibm_marrakesh for nine set of qubits. (a) shows the logical errors and (b) shows the postselection results. This device features tunable couplers that significantly reduce the crosstalk. Nevertheless, accumulation of dephasing errors over time (up to $\approx 55\leavevmode\nobreak\ \mu$s) lead to the appearance of logical errors. The same LDD sequence we used in the main text successfully suppresses these errors, as shown in (a). Additionally, LDD with postselection outperforms postselection alone, as shown in (b).
Thus, the unencoding circuit, i.e., the circuit mapping every logical and stabilizer operator from the $[[4,2,2]]$ code to the corresponding operators of the trivial code, is
$V^{\dagger}=\text{CNOT}_{34}H_{4}H_{3}\text{CNOT}_{24}\text{CNOT}_{42}\text{CNOT}_{21}\text{CNOT}_{14},$ (11)
where $H_{i}$ is the Hadamard on the $i$th qubit.
After unencoding, we obtain the trivial code with $\mathcal{S}_{0}=\langle Z_{3},Z_{4}\rangle$ and $\mathcal{L}_{0}=\langle X_{1},X_{2},Z_{1},Z_{2}\rangle$, that divides the code into $k=2$ data qubits and $n-k=2$ ancillas.
In the trivial code, error correction is as follows: apply $IIXI$ when $IIZI=-1$ and $IIIX$ when $IIIZ=-1$. Applying $V^{\dagger}$ to these two recovery operators we find that to correct an undecoupled error in the original $[[4,2,2]]$ code, one should apply $IIIZ$ whenever $XXXX=-1$ and $XXXI$ (or $IIIX$, since they are related by a stabilizer element) when $ZZZZ=-1$.
Appendix D Issues involving the implementation of logical state tomography using the $[[4,2,2]]$ code
A general two-qubit state requires information about $15$ expectation values which requires $9$ independent measurement settings (achieved by measuring each qubit in three complementary bases; e.g., $X$, $Y$, and $Z$[103]. This holds for logical tomography of a logical two-qubit state as well. There are a few ways in which a logical observable can be measured. First, one can use stabilizer-measurement-like circuits where the logical operator is measured by executing a circuit composed of CNOT gates targeting an ancilla qubit that is measured to learn the logical measurement outcome. This method can further be paired with a round of syndrome extraction using additional ancilla qubits to learn whether the state was in the code space to begin with. This combination of measurements allows us to invoke the code’s protection while simultaneously performing a measurement. The downside of this method is the additional overhead in CNOTs and ancilla qubits needed to perform the measurement protocol. In our case, this would require a substantial overhead in SWAP gates as well since the IBM QPU’s heavy-hex lattice does not pair naturally with the $[[4,2,2]]$ code. Using this logical measurement method would inevitably introduce more errors and reduce the accuracy of logical state tomography.
An alternative, less costly approach for performing logical measurements is to directly measure all data qubits of the code in lieu of introducing ancillas and additional CNOTs. However, this method is incompatible with extracting information about the entire stabilizer generator set while also performing the logical measurement. As a result, we would not be able to know with certainty that the system state was in the code space at the time of measurement. There is some nuance to this approach, which does allow us to learn information about stabilizers which commute qubit-wise (i.e., they share the same Pauli operator or $I$ on the same qubit) with the logical operator being measured. For example, we can simultaneously measure $\overline{X}_{1}=XIIX$, $\overline{X}_{2}=IIXX$, $\overline{X_{1}X_{2}}=XIXI$ and the stabilizer $S_{1}=XXXX$ since these operators all qubit-wise commute. Performing the measurement of each data qubit in the $X$ basis gives us information about each of these logical observables, as well as information about whether or not the system was in the logical code space with respect to the stabilizer generator $S_{1}$. Unfortunately, this symmetry does not hold for the operators $\overline{Y}_{1}=XIZY$, $\overline{Y}_{2}=ZIXY$, and $S_{2}=YYYY$ since these operators do not all commute qubit-wise. This means that we cannot simultaneously learn information about these logical observables and information about whether or not the system was in the code space with respect to the stabilizer $S_{2}$ in the same shot, reducing our total information about the system. This would lead to a less precise estimation of the logical density matrix, as we will count more experimental shots involving the system outside of the logical state space in our estimation of the logical expectation value.
For these reasons, we did not use logical state tomography in this work.
Appendix E Logical dynamical decoupling with tunable couplers
Here we present additional experiments on ibm_marrakesh, which features tunable couplers [104] unlike the always-on $ZZ$ interaction in ibm_kyiv. Consequently, for this set of experiments, crosstalk is significantly reduced compared to the other set of results ($\lesssim 5$ kHz vs tens of kHz). Nevertheless, a combination of dephasing and residual crosstalk accumulates, leading to logical errors. Figure 8 illustrates the performance of LDD sequences on this QPU, averaged over 9 different qubit sets. Fig. 8(a) [equivalent to Fig. 3(b)] shows that without DD, logical errors remain unsuppressed. However, both logical DD sequences we employ successfully suppress these errors. Fig. 8(b) [equivalent to Fig. 5(a)] demonstrates that using only postselection, the fidelity averages to $\mathcal{F}\approx 56.23\%$, whereas combining postselection with LDD increases it to $\mathcal{F}\approx 91.73\%$.
Optimizing LDD for specific types of tunable couplers to maximize the interplay between DD and couplers is an avenue for future work.

### v2

@@
 Here, we propose and implement a method that leverages dynamical decoupling (DD) to drastically suppress logical errors.
-The key to achieving this is to use the logical operators of the QEC code as DD pulses, which we refer to as logical dynamical decoupling (LDD).
-The resulting hybrid QEC-LDD strategy is in principle capable of handling arbitrary weight errors.
-We test QEC-LDD using IBM transmon devices and the $[[4,2,2]]$ code, demonstrating performance that significantly exceeds the capabilities of using either this code or DD in isolation.
+The key to achieving this is to use the normalizer elements of the QEC code as DD pulses, which we refer to as normalizer dynamical decoupling (NDD).
+The resulting hybrid QEC-NDD strategy is in principle capable of handling arbitrary weight errors.
+We test QEC-NDD using IBM transmon devices and the $[[4,2,2]]$ code, demonstrating performance that significantly exceeds the capabilities of using either this code or DD in isolation.
 We present a method that allows for the detection of logical errors affecting logically encoded Bell states, which, in this case, arise primarily from crosstalk among physical qubits.
 Building on this, we experimentally demonstrate high-fidelity entangled logical qubits.
+The fidelities we achieve are beyond-breakeven, i.e., they significantly exceed the corresponding fidelities of unprotected entangled qubits in the same setting.
+I Introduction
 Quantum error correction (QEC) [1, 2, 3, 4] is fundamental to the realization of fault-tolerant quantum computation [5, 6, 7, 8], ensuring the preservation of quantum information undergoing error processes during computation and storage [9, 10].
-Numerous successful experimental demonstrations of QEC across various platforms have been reported over the years [11, 12, 13, 14, 15], with the scale and pace accelerating recently towards genuine fault tolerance [16, 17, 18, 19, 18, 20, 21, 22, 23, 24].
-Fault-tolerant quantum computing will require the stability of logical qubits over the potentially hours-long timescales of quantum algorithms solving utility-scale problems [25, 26, 27].
-Above threshold, over such timescales low-weight physical errors may transform into logical errors of weight higher than any deployed fixed-distance QEC code can handle.
-This can become a concern even below threshold since long-range spatial and temporal correlations may develop that violate the assumptions underlying fault tolerance theory [28, 29, 30, 31, 32].
+Numerous successful experimental demonstrations of QEC across various platforms have been reported over the years [11, 12, 13, 14, 15], with the scale and pace accelerating recently towards genuine fault tolerance [16, 17, 18, 19, 20, 21, 22, 23, 24].
+Fault-tolerant quantum computing will require the stability of logical qubits on the long timescales of quantum algorithms that solve utility-scale problems [25, 26, 27].
+Above threshold, on such timescales, low-weight physical errors may transform into logical errors, i.e., errors of weight higher than a deployed fixed-distance QEC code can handle.
+This can become a concern even below threshold, as long-range spatial and temporal correlations may develop that violate the assumptions underlying fault tolerance theory [28, 29, 30, 31, 32].
 Conventionally, suppressing higher-weight errors requires increasing the QEC code distance, e.g., by means of code concatenation [33, 34], or increasingly larger codes such as is done with surface codes [35, 36], color codes [37], or quantum low-density parity-check codes [38].
@@
 Here, we propose and demonstrate a low-cost method that combines a fixed-distance QEC or quantum error detection (QED) code with dynamical decoupling (DD) [39, 40, 41, 42].
-This hybrid method, which we call QEC-LDD, can handle arbitrary-weight errors.
-Here, LDD stands for ‘logical dynamical decoupling’; the decoupling sequence is implemented using the logical operators of the QEC code as pulses.
-We prove a corresponding QEC-LDD theorem which, informally, posits that the most general set of errors acting on an $[[n,k,d]]$ code, namely errors with weights up to $n$, can all either be suppressed using LDD or corrected using the code, for any $d\geq 1$.
-This seemingly paradoxical result is enabled by virtue of the fact that the LDD sequence suppresses all logical errors as well as some physical errors, thus leaving a much smaller subset of errors for the code to handle, which turn out to all be correctable.
-Notably, these are not the errors the code was originally designed to correct; indeed, the result holds in particular also for purely error detecting codes.
-The QEC-LDD theorem is a formal result, in the sense that QEC-LDD, if implemented literally as described in the theorem, is not expected to outperform a pure DD protocol (absent any QEC) on $k$ physical qubits.
-Nevertheless, the QEC-LDD theorem serves as a useful framework for less ambitious protocols that do not attempt to remove all errors but instead focus on the more physically relevant subset.
-We discuss and illustrate this conclusion in detail in this work, in the context of a transmon-based demonstration of entangled logical qubits.
-Standard, ‘physical’ DD, where DD pulses act not at the logical level but at the physical qubit level, has recently shown great progress, improving the fidelity of storing quantum states [43, 44, 45, 46, 47, 48, 49], circuits [50, 51, 52], and even the performance of entire algorithms [53, 54, 55, 56].
+This hybrid method, which we call QEC-NDD, can handle arbitrary-weight errors.
+Here, NDD stands for “normalizer dynamical decoupling”; the decoupling sequence is implemented using the normalizer elements of the QEC code as pulses.
+Standard, “physical” DD, where DD pulses act not at the logical level but at the physical qubit level, has recently shown great progress, improving the fidelity of storing quantum states [43, 44, 45, 46, 47, 48, 49], circuits [50, 51, 52], and even the performance of entire algorithms [53, 54, 55, 56].
 Furthermore, physical DD can be seamlessly combined with fault-tolerant QEC [57, 58] and several recent QEC experiments have used physical DD profitably [18, 21, 19, 59, 20, 60].
 However, such pulse sequences can introduce additional errors due to control imperfections in the pulses and due to crosstalk, potentially overshadowing the benefits of DD.
-To address this challenge, we design our LDD sequences to be robust against such control errors [61, 62] and crosstalk [63, 64, 65, 66, 67], ensuring that the advantages of logical error suppression are not compromised.
-This robust design allows our LDD implementation to enhance the protection of the code space as intended, effectively targeting both logical and control errors.
+To address this challenge, we design our NDD sequences to be robust against such control errors [61, 62] and crosstalk [63, 64, 65, 66, 67], ensuring that the advantages of logical error suppression are not compromised.
+This robust design allows our NDD implementation to enhance the protection of the code space as intended, effectively targeting both logical and control errors.
 We demonstrate the practical utility and advantages of this approach using $127$-qubit IBM quantum processors.
 The workhorse in our demonstrations is the $[[4,2,2]]$ quantum error detection code, whose two logical qubits we use to prepare logically encoded Bell states.
-To demonstrate the effectiveness of LDD in suppressing logical errors, we first need a method to detect such errors.
+To demonstrate the effectiveness of NDD in suppressing logical errors, we first need a method to detect such errors.
 However, this code is constrained by its low distance and is incapable of detecting prevalent logical errors, including $ZZ$ crosstalk errors.
 To overcome this limitation, we design experiments in which, through the use of logically encoded Bell states, we nevertheless unequivocally detect the occurrence of logical errors for a known input logical state.
-This is then followed by the implementation of various versions of LDD, each employing different subsets of logical operators to construct the sequence.
+This is then followed by the implementation of various versions of NDD, each employing different subsets of normalizer elements to construct the sequence.
 These implementations effectively demonstrate significant suppression of logical errors and substantial improvement in the fidelity of the code space.
 The structure of this paper is as follows.
-In Section I we provide pertinent background on the error model and DD from the perspective of group averaging.
-We discuss standard Pauli-group DD as well as DD implemented using the stabilizers and or logical operators of a stabilizer code.
-In Section II we state and prove the QEC-LDD Theorem.
-We also illustrate the theorem using the example of the $[[4,2,2]]$ code.
-We then turn to our experimental demonstration of high-fidelity logical Bell states in Section III.
-This section describes our experimental design and methodology, evidence that both physical DD and error detection with postselection improve logical Bell state fidelity, and finally the evidence that the hybrid QEC-LDD strategy is significantly better.
-We conclude in Section IV and provide supporting technical details in the Appendix.
-I Background
-I.1 Quantum codes and logical errors
-An $[[n,k,d]]$ code encodes $k$ logical qubits into $n$ physical qubits with distance $d$ [68].
-The weight of an error is the number of physical qubits it affects simultaneously.
-The highest-weight errors an $[[n,k,d]]$ code can correct or detect are $\lfloor\frac{d-1}{2}\rfloor$ and $d-1$, respectively.
-A $d=2$ code is a pure quantum error detection (QED) code, and a distance-$d$ QEC code can always be used as a QED code for errors of weight $<d$.
-The code distance is the minimum number of physical qubits that must experience an error to cause an undetectable logical error, i.e., an error forming a logical operation inside the code space.
-These logical errors can either be inherently present or result from the accumulation of lower-weight errors over time.
-For example, in the context of superconducting qubits with fixed-frequency couplers, a prevalent challenge is $ZZ$ crosstalk [69, 70], which inherently introduces weight-two errors that can present as logical errors for distance-$2$ codes.
-I.2 Logical dynamical decoupling
-Let $\mathcal{P}_{n}=\{\pm P_{j},\pm iP_{j}\}_{j=0}^{4^{n}-1}\subset\mathsf{U}(2^{n})$ be the full Pauli group with $4^{n+1}$ elements generated by $\{I,X,Y,Z\}^{\otimes n}$ and the phase factors {±$\{\pm 1,\pm i\}$.
-Let $\tilde{\mathcal{P}}_{n}=\mathcal{P}_{n}/\{\pm 1,\pm i\}\subset\mathsf{U}(2^{n})/\mathsf{U}(1)$ be the Abelian factor group with $4^{n}$ elements (here $\mathsf{U}(m)$ is the group of unitary operators on $\mathbb{C}^{m}$).
-Consider the most general ‘total decoherence’ system-bath interaction for $n$ qubits:
-$H_{SB}=\sum_{j=0}^{4^{n}-1}\alpha_{j}P_{j}\otimes B_{j}.$ (1)
-We let $P_{0}=I^{\otimes n}$, the $n$-qubit identity operator. The pure-bath Hamiltonian $P_{0}\otimes H_{B}$ is harmless.
-In its simplest form, accounting only for first-order decoupling and ignoring pulse errors, DD theory can be understood as group symmetrization [71, 72].
-Choose a decoupling group $\tilde{\mathcal{G}}$ that consists of a set of unitary transformations (pulses) $g_{j}$ acting purely on the system: $\tilde{\mathcal{G}}=\left\{g_{j}\right\}_{j=0}^{\left|\tilde{\mathcal{G}}\right|-1}$, where $g_{0}=P_{0}$.
-We assume that each such pulse is a unitary operator defined up to a global phase, i.e.
-$\tilde{\mathcal{G}}\subset\mathsf{U}(2^{n})/\mathsf{U}(1)$; however, to simplify the notation we write, e.g., $g_{0}=P_{0}$ instead of $g_{0}=[P_{0}]$ (the equivalence class of $P_{0}$ up to a global phase).
-Similarly to $\mathcal{P}_{n}$ and $\tilde{\mathcal{P}}_{n}$, we assume that $\tilde{\mathcal{G}}$ can be written as $\tilde{\mathcal{G}}=\mathcal{G}/\{\pm 1,\pm i\}$ for some group $\mathcal{G}\subset\mathsf{U}(2^{n})$.
-In particular, we assume that the pulses are instantaneous and cycle over all elements of the group $\tilde{\mathcal{G}}$.
-This yields the DD pulse sequence:
-$U(T)=\prod_{j=0}^{\left|\tilde{\mathcal{G}}\right|-1}g_{j}^{\dagger}f_{\tau}g_{j}=e^{-iT(\langle H_{SB}\rangle_{\mathcal{G}}+P_{0}\otimes H_{B})}+O(T^{2}).$ (2)
-Here, $\tau$ is the pulse interval, $T=|\tilde{\mathcal{G}}|\tau$ is the total duration of one repetition of the sequence, $f_{\tau}=e^{-i\tau H_{SB}}$ is the free-evolution unitary, and $\langle H_{SB}\rangle_{\mathcal{G}}$ is the group-symmetrized system-bath Hamiltonian, where
-$\langle A\rangle_{\mathcal{G}}\equiv\frac{1}{|\mathcal{G}|}\sum_{g\in\mathcal{G}}g^{\dagger}Ag=\frac{1}{|\tilde{\mathcal{G}}|}\sum_{g\in\tilde{\mathcal{G}}}g^{\dagger}Ag$ (3)
+In Section II we provide pertinent background on the error model and DD from the perspective of group averaging.
+We discuss standard Pauli group DD, as well as encoded DD, i.e., DD implemented using the stabilizers and/or logical operators of a stabilizer code.
+In Section III we describe our experimental design and methodology.
+We then turn to our experimental demonstration of high-fidelity logical Bell states in Section IV.
+This section describes our experimental design and methodology, evidence that both physical DD and error detection with postselection improve logical Bell state fidelity, and finally, the evidence that the hybrid QEC-NDD strategy significantly outperforms both standalone dynamical decoupling and error detection.
+We conclude in Section V and provide supporting technical details in the Appendix.
+II Background
+In this section, we provide all the theoretical dynamical decoupling background necessary in order to understand the experimental results presented in Section III.
+In particular, after defining our Hamiltonian error model, we explain how different flavors of encoded DD, i.e., DD implemented using the stabilizers and/or logical operators of a QEC code, suppress the different types of errors that arise in our error model.
+II.1 Error model
+Let $\mathsf{U}(m)$ denote the group of unitary operators on $\mathbb{C}^{m}$.
+Let $\mathcal{P}_{n}=\{\pm P_{j},\pm iP_{j}\}_{j=0}^{4^{n}-1}\subset\mathsf{U}(2^{n})$ be the full Pauli group with $4^{n+1}$ elements generated by $\{I,X,Y,Z\}^{\otimes n}$ and the phase factors {±$\{\pm 1,\pm i\}$, where $X\equiv\sigma^{x}$ and likewise for $Y$ and $Z$.
+There is a natural projection (quotient homomorphism) $\pi:\mathcal{P}_{n}\to\tilde{\mathcal{P}}_{n}$ which takes the Pauli operator and ignores its global phase, yielding a “phase-stripped” Pauli.
+Correspondingly, let $\tilde{\mathcal{P}}_{n}=\mathcal{P}_{n}/\{\pm 1,\pm i\}\subset\mathsf{U}(2^{n})/\mathsf{U}(1)$ be the Abelian factor group with $4^{n}$ elements, also known as the projective Pauli group.
+Consider the most general “total decoherence” system-bath interaction for $n$ qubits:
+$H_{SB}=\sum_{{j=1}}^{4^{n}-1}\alpha_{j}P_{j}\otimes B_{j}.$ (1)
+In writing this expression we exclude the bath Hamiltonian $I_{S}\otimes B_{0}$ (where $B_{0}=H_{B}$) and select exactly one representative $P_{j}$ from each of the $4^{n}$ equivalence classes $\tilde{\mathcal{P}}_{n}$ such that (s.t.) $P_{0}=I_{S}$ (we use the notation $I_{S}$ to denote the identity operator on the entire system Hilbert space; we use $I$ for a generic identity operator).
+There is still some freedom in how we order the elements of $\tilde{\mathcal{P}}_{n}$ and which phases we choose.
+We use this freedom later to organize $\{P_{j}\}_{j=1}^{4^{n}-1}$ in the context of $[[n,k,d]]$ codes.
+Our model for $H_{SB}$ is general enough to contain system-only error terms, since any of the bath operators $B_{j}$ may be set equal to the bath identity operator $I_{B}$.
+Henceforth, when unambiguous, we use $I$ to denote the identity operator regardless of the space on which it acts or its dimension.
+The joint system-bath free-evolution unitary is $f_{\tau}=e^{-i\tau(H_{SB}+I\otimes H_{B})}$.
+The adjective “free” refers to the fact that we are leaving out any system-only control Hamiltonian; such control will enter later through DD or QEC.
+The goal of the latter is, respectively, to suppress or correct the deleterious effects caused by $H_{SB}$.
+II.2 Dynamical decoupling and quantum error correction
+In its simplest form, accounting only for first-order decoupling and ignoring pulse errors, DD theory can be understood as group symmetrization [68, 69, 70].
+Consider a discrete group $\mathcal{G}\subset\mathsf{U}(2^{n})$ with elements $\{g_{j}\}_{j=0}^{|\mathcal{G}|-1}$ representing unitary transformations $g_{j}$ acting purely on the system.
+We set $g_{0}\equiv g_{|\mathcal{G}|}\equiv I_{S}$ and refer to the corresponding phase-stripped group $\tilde{\mathcal{G}}=\pi(\mathcal{G})=\left\{\tilde{g}_{j}\right\}_{j=0}^{K-1}\subset\mathsf{U}(2^{n})/\mathsf{U}(1)$ as the “decoupling group”, where $K\equiv|\tilde{\mathcal{G}}|=|\mathcal{G}|/4$.
+This yields the DD cycle
+$U(T)$ $=\prod_{j=0}^{K-1}\tilde{g}_{j}^{\dagger}f_{\tau}\tilde{g}_{j}=e^{-iT(\langle H_{SB}\rangle_{\mathcal{G}}+I\otimes H_{B})}+O(T^{2})$ (2)
+$=p_{K}f_{\tau}p_{K-1}f_{\tau}\cdots f_{\tau}p_{1}f_{\tau}.$
+Here, $\{p_{j}=\tilde{g}_{j}\tilde{g}_{j-1}^{\dagger}\}_{j=1}^{K}$ are the pulses of the corresponding DD pulse sequence, $\tau$ is the pulse interval, and $T=K\tau$ is the total duration of one repetition of the sequence.
+Note that, equivalently, $\tilde{g}_{j-1}=p_{j-1}\cdots p_{0}$, where $p_{0}\equiv I_{S}$.
+$\langle H_{SB}\rangle_{\mathcal{G}}$ is the group-symmetrized system-bath Hamiltonian, where
+$\langle A\rangle_{\mathcal{G}}\equiv\frac{1}{|\mathcal{G}|}\sum_{g\in\mathcal{G}}g^{\dagger}Ag=\frac{1}{K}\sum_{\tilde{g}\in\tilde{\mathcal{G}}}\tilde{g}^{\dagger}A\tilde{g}$ (3)
 is the projection of the arbitrary bounded system-bath operator $A$ onto the subalgebra of operators that commute with every element of $\mathcal{G}$ (i.e., its commutant).
-We say that $\tilde{\mathcal{G}}$ decouples $H_{SB}$ (to first order) if $\langle H_{SB}\rangle_{\mathcal{G}}=cI$ where $c$ is a constant, including zero.
+$\langle A\rangle_{\mathcal{G}}$ is also called the group average of $A$ with respect to $\mathcal{G}$.
+Definition 1 (Decoupling).
+Let $A$ be an arbitrary operator acting on the joint system-bath Hilbert space.
+$\tilde{\mathcal{G}}$ decouples $A$ (to first order) if $\langle A\rangle_{\mathcal{G}}=cI\otimes B$ where $c$ is a constant, including zero, and $B$ is an arbitrary bath operator.
 For example, $\tilde{\mathcal{G}}=\tilde{\mathcal{P}}_{n}$ decouples an arbitrary $n$-qubit system-bath Hamiltonian since then $\langle H_{SB}\rangle_{\mathcal{G}}=0$, albeit at a sequence time cost of $T=4^{n}\tau$ [40].
-To illustrate this, consider $n=1$; the most general system-bath Hamiltonian of a single qubit is $H_{SB}=\sum_{i=0}^{3}\alpha_{i}P_{i}\otimes B_{i}$ with $P_{i}\in\{I,X,Y,Z\}$.
-If $\tilde{\mathcal{G}}=\{I,X,Y,Z\}$, then $U(T)$ simplifies into the well-known XY4 sequence $U(4\tau)=Yf_{\tau}Xf_{\tau}Yf_{\tau}Xf_{\tau}$ [73], and indeed, $\langle H_{SB}\rangle_{\mathcal{G}}=0$.
-Higher-order decoupling sequences are known that achieve suppression up to arbitrary order $q$ [i.e., leaving an $O(T^{q+1})$ unsuppressed error term in $U(T)$[74, 75, 76, 77, 78], but for simplicity, in this work we restrict our attention to first-order sequences.
-Note that a common misconception is that DD is not effective against purely Markovian noise [79]; however this is not the case, essentially since even in the Markovian limit the bath can have a nonzero correlation time [80, 81, 82, 83].
-[…]
-For a code $\mathcal{C}\subset(\mathbb{C}^{2})^{\otimes n}$, the stabilizer group $\mathcal{S}$ is uniquely defined as
-$\mathcal{S}=\{P\in\mathcal{P}_{n}:\forall\ket{\psi}\in\mathcal{C},\;P\ket{\psi}=\ket{\psi}\}.$ (4)
-Its centralizer is also uniquely defined.
-However, the group of canonical logical operators is not unique: one can choose arbitrary stabilizers $S_{j}^{X},S_{j}^{Z}\in\mathcal{S}$ ($j=1,\dots,k$) and multiply the generators $\bar{X}_{j},\bar{Z}_{j}$ by those to obtain other choices of the group $\mathcal{L}$.
-It is always possible to write the system-bath interaction [Eq. 1] as
-$H_{SB}=H_{SB}^{\mathcal{S}}+H_{SB}^{\mathcal{L}}+H_{SB}^{\mathcal{E}},$ (5)
-grouping together terms with similar action on the code space $\mathcal{C}$.
-Altogether, this expression has $4^{n}$ linearly independent terms corresponding to the elements of $\tilde{\mathcal{P}}_{n}$.
-Specifically, $H_{SB}^{\mathcal{S}}$ collects all the distinct $2^{n-k}$ Pauli operators $P_{i}\in\mathcal{S}$ (including $P_{0}=I$); such terms act as the identity operator on the code space and, therefore, are harmless.
-$H_{SB}^{\mathcal{L}}$ collects all $2^{n-k}(4^{k}-1)$ Pauli operators in $(\mathcal{S}\mathcal{L}\setminus\mathcal{S})/\{\pm 1,\pm i\}$; these terms leave the code space (and each of the other syndrome spaces) invariant but cause harmful logical errors, since by construction, the code cannot detect such errors.
-Finally, $H_{SB}^{\mathcal{E}}$ collects all remaining $2^{n+k}(2^{n-k}-1)$ terms in $\tilde{\mathcal{P}}_{n}\setminus(\mathcal{S}\mathcal{L}/\{\pm 1,\pm i\})$, which correspond to potentially detectable errors: given the value of the syndrome one can either detect all of them or correct some of them.
-A key observation is that the decoupling group is arbitrary and may, in particular, be formed from the stabilizer group $\mathcal{S}$ and/or the group of canonical logical operators $\mathcal{L}$ of a stabilizer code $\mathcal{C}$, or even a subset of either one [87, 72, 88, 89, 58, 90].
-One approach for selectively combining DD with QEC, which leverages this flexibility, is to use DD only to suppress the errors between or within the $2^{n-k}$ syndrome spaces, including the code space $\mathcal{C}$ (i.e., errors in $(\mathcal{S}\mathcal{L}\setminus\mathcal{S})/\{\pm 1,\pm i\}$).
-Errors causing transitions between the syndrome spaces (i.e., elements of $\tilde{\mathcal{P}}_{n}\setminus(\mathcal{S}\mathcal{L}/\{\pm 1,\pm i\})$) are the ones the code aims to detect or correct.
-The former are decoupled by $\mathcal{G}=\mathcal{L}$ while the latter are decoupled by $\mathcal{G}=\mathcal{S}$.
-For any choice of $\mathcal{G}$, the decoupling condition is that there exists at least one element of $\mathcal{G}$ that anticommutes with a given error.
-Based on these properties, it was shown in Ref.
-[58] that combining the full set of $n+k$ stabilizer and canonical logical operator generators to generate the DD group $\mathcal{G}=\mathcal{S}\times\mathcal{L}$, optimally achieves the goal of decoupling the errors between or within the $2^{n-k}$ syndrome spaces, in the sense of minimizing the length of the DD pulse sequence.
-This is known as ‘stabilizer-logical DD’ (SLDD) and incurs a single sequence cost of $T=2^{n+k}\tau$, instead of $4^{n}\tau$ in the case of full Pauli group decoupling.
+To illustrate this, consider $n=1$; the most general system-bath Hamiltonian of a single qubit is $H_{SB}=\sum_{{i=1}}^{3}\alpha_{i}P_{i}\otimes B_{i}$ with $P_{i}\in\mathcal{P}_{1}\setminus I$.
+If $\tilde{\mathcal{G}}={\tilde{\mathcal{P}}_{1}}=\langle X,Z\rangle=\{I,X,Y,Z\}$, then $U(T)$ simplifies into the well-known (universal) XY4 sequence $U(4\tau)=Yf_{\tau}Xf_{\tau}Yf_{\tau}Xf_{\tau}$ [71], and indeed, $\langle H_{SB}\rangle_{\mathcal{G}}=0$.
+The XX sequence is $U(2\tau)=Xf_{\tau}Xf_{\tau}$, and decouples $Y$ and $Z$, but not $X$ errors.
+The XY4 sequence can be seen as a concatenation of the XX and ZZ sequences, with the latter defined similarly to XX [72].
+Note that a common misconception is that DD is not effective against purely Markovian noise; however, this is not the case, essentially since even in the Markovian limit, the bath can have a nonzero correlation time [73, 74, 75, 76].
+In general, decoupling using subgroups of $\tilde{\mathcal{P}}_{n}$ will eliminate parts of $H_{SB}$, presenting an opportunity to selectively combine DD with QEC [77, 57].
+See Appendix A for pertinent background on QEC codes.
+Let $\mathcal{S}$ be a stabilizer group specifying an $[[n,k,d]]$ stabilizer code.
+The group of logical operators is $\mathcal{N}(\mathcal{S})/\mathcal{S}$, where $\mathcal{N}(\mathcal{S})$ is the normalizer of $\mathcal{S}$ in $\mathcal{P}_{n}$ (which here coincides with the centralizer, i.e., the group of Pauli operators commuting with all elements of $\mathcal{S}$).
+We can pick a canonical set of generators of the stabilizer group and a canonical group of logical operators $\mathcal{L}$ [^1] That is, any transversal for $\mathcal{S}$ in $\mathcal{N}(\mathcal{S})$ satisfying $\pi^{-1}(\pi(\mathcal{L}))=\mathcal{L}$ can be picked as $\mathcal{L}$, i.e., any subgroup of $\mathcal{N}(\mathcal{S})$ with exactly one representative from each coset $L\mathcal{S}\subset\mathcal{N}(\mathcal{S})$ s.t.
+$iI\in\mathcal{L}$.
+Any such $\mathcal{L}$ is isomorphic to $\mathcal{P}_{k}$ and can be thought of as a group of Paulis acting on logical qubits.
+We can use the remaining freedom mentioned above to reorder $\{P_{j}\}_{j=0}^{4^{n}-1}$ and pick the phases of those operators to ensure that
+  • $P_{0}=I_{S}$
+  • $\mathcal{S}_{\star}=\{P_{1},\dots,P_{2^{n-k}-1}\}=\mathcal{S}\setminus\{I_{S}\}$ (these are non-identity stabilizers acting trivially on the code space)
+  • =$\mathcal{L}_{\star}=\{P_{2^{n-k}},\dots,P_{2^{n+k}-1}\}\subset\mathcal{N}(\mathcal{S})\setminus\pi^{-1}(\tilde{\mathcal{S}})$ (these manifest as undetectable logical errors when acting on the code space)
+  • $\mathcal{D}_{\star}=\{P_{2^{n+k}},\dots,P_{4^{n}-1}\}\subset\mathcal{P}_{n}\setminus\mathcal{N}(\mathcal{S})$ (these are detectable errors).
+We chose the phases of the representatives of $\tilde{\mathcal{S}}$ to ensure they are actually in $\mathcal{S}$ (and not different from elements of $\mathcal{S}$ by a phase).
+This represents $\{P_{j}\}_{j=0}^{4^{n}-1}$ as a disjoint union
+$\{P_{j}\}_{j=0}^{4^{n}-1}=\{I_{S}\}\dot{\cup}\mathcal{S}_{\star}\dot{\cup}\mathcal{L}_{\star}\dot{\cup}\mathcal{D}_{\star}.$ (4)
+The system Hilbert space $\mathcal{H}$ can be split as
+$\mathcal{H}=\mathcal{H}_{\text{stab}}\otimes\mathcal{H}_{\text{log}},$ (5)
+where $\mathcal{S}$ acts on the first component and $\mathcal{L}$ acts on the second, and where $\dim(\mathcal{H}_{\text{stab}})=2^{n-k}$ and $\dim(\mathcal{H}_{\text{log}})=2^{k}$, corresponding to the $n-k$ (virtual) syndrome qubits and $k$ logical qubits, respectively.
+With respect to this virtual tensor product decomposition induced by the syndrome measurement [78], every Pauli operator $E\in\mathcal{S}_{\star}$, $\mathcal{L}_{\star}$, or $\mathcal{D}_{\star}$ can be written as:
+  • $E=P\otimes I\in\mathcal{S}_{\star}$ with $P\in\mathcal{P}_{n-k}$ (a stabilizer with trivial action on the code space)
+  • $E=P\otimes Q\in\mathcal{L}_{\star}$ with $[P,\mathcal{S}_{\star}]=0$ (a logical operator manifesting as an error if it appears in $H_{SB}$).
+  • $E=P\otimes Q\in\mathcal{D}_{\star}$ with $\{P,S_{i}\}=0$ for one or more $S_{i}\in\mathcal{S}_{\star}$ (a detectable error).
+Correspondingly, it is always possible to write the system-bath interaction [Eq. 1] as
+$H_{SB}$ $=H_{SB}^{\mathcal{S}_{\star}}+H_{SB}^{\mathcal{L}_{\star}}+H_{SB}^{\mathcal{D}_{\star}}$ (6a)
+$H_{SB}^{\mathcal{S}_{\star}}\!\!=\sum_{i=1}^{2^{n-k}-1}S_{i}\otimes B^{\text{stab}}_{i}\ ,\quad S_{i}\in\mathcal{S}_{\star}$ (6b)
+$H_{SB}^{\mathcal{L}_{\star}}=\sum_{i=1}^{2^{n-k}(4^{k}-1)}L_{i}\otimes B^{\text{log}}_{i}\ ,\quad L_{i}\in\mathcal{L}_{\star}$ (6c)
+$H_{SB}^{\mathcal{D}_{\star}}=\sum_{i=1}^{4^{n}-2^{n+k}}D_{i}\otimes B^{\text{det}}_{i}\ ,\quad D_{i}\in\mathcal{D}_{\star},$ (6d)
+where $\{B^{\text{stab}}_{i},B^{\text{log}}_{i},B^{\text{det}}_{i}\}$ are bath operators with units of energy.
+Eq.
+6 groups terms with similar action on the code space $\mathcal{C}$.
+In total, $H_{SB}+I_{S}\otimes H_{B}$ has $4^{n}$ linearly independent terms corresponding to the elements of $\tilde{\mathcal{P}}_{n}$.
+Specifically, $H_{SB}^{\mathcal{S}_{\star}}+I_{S}\otimes H_{B}$ collects all $2^{n-k}$ distinct terms in $\{I_{S}\}\dot{\cup}\mathcal{S}_{\star}$ that act as the identity operator on the code space and, therefore, are harmless.
+$H_{SB}^{\mathcal{L}_{\star}}$ collects all $2^{n-k}(4^{k}-1)$ Pauli operators in $\mathcal{L}_{\star}$; these terms are non-trivial logical operators that leave the code space (and each of the other syndrome spaces) invariant.
+Finally, $H_{SB}^{\mathcal{D}_{\star}}$ collects all remaining $4^{n}-2^{n-k}-2^{n-k}(4^{k}-1)=4^{n}-2^{n+k}$ terms in $\mathcal{D}_{\star}$, which correspond to detectable and potentially correctable) errors.
+We can now establish when an error $E\in\mathcal{P}_{n}$ is decoupled.
+Lemma 1.
+Let $\mathcal{H}=(\mathbb{C}^{2})^{\otimes n}$ be the Hilbert space of $n$ qubits.
+Then all elements of $\mathcal{P}_{n}$ are unitary operators on $\mathcal{H}$.
+Consider a term in $H_{SB}$ corresponding to some $E\in\mathcal{P}_{n}$.
+It is decoupled by $\tilde{\mathcal{G}}$ in the sense of 1 if and only if there exists $g\in\mathcal{G}$ anticommuting with $E$.
+Proof.
+If there exists $g\in\mathcal{G}$ that anticommutes with $E$ then $g^{\dagger}Eg=-E$ and
+$\langle E\rangle_{\mathcal{G}}=\frac{1}{\left|\mathcal{G}\right|}\sum_{h\in\mathcal{G}}h^{\dagger}Eh=-\frac{1}{\left|\mathcal{G}\right|}\sum_{h\in\mathcal{G}}(gh)^{\dagger}E{gh}=-\langle E\rangle_{\mathcal{G}},$ (7)
+so the group average $\langle E\rangle_{\mathcal{G}}=0$.
+If there is no such $g$, then all elements of $\mathcal{G}$ commute with $E$ (because every two elements of $\mathcal{P}_{n}$ either commute or anticommute).
+Then $g^{\dagger}Eg=E$ for all $g\in\mathcal{G}$, hence $\langle E\rangle_{\mathcal{G}}=E$ (i.e., the term is undecoupled).
+∎
+II.3 Encoded dynamical decoupling
+A key observation is that the decoupling group is arbitrary [70, 69] and may, in particular, be formed from the stabilizer group $\mathcal{S}$ and/or the group of logical operators $\mathcal{L}$ of a stabilizer code $\mathcal{C}$ [58].
+We call these different choices “encoded dynamical decoupling,” as they all involve the use of encoded operations (relative to a QEC code) as DD pulses, instead of physical DD pulses.
+We next provide several examples of encoded DD groups.
+When using a QEC code, the approach of using the full Pauli group $\mathcal{P}_{n}$ is excessive, since there is no need to decouple the terms in $H_{SB}^{\mathcal{S}_{\star}}$; this observation motivated Ref.
+[58] to introduce SLDD, which achieves a reduction from a DD group order of $4^{n}$ to $2^{n+k}$.
 The savings are substantial when $k\ll n$.
-Adding QEC syndrome measurements and recovery operations based on syndrome measurements then detects the remaining errors between syndrome spaces that DD did not suppress [i.e., the $O(T^{q+1})$ terms in the case of a $q$’th-order DD sequence] or successfully corrects some of them, resulting in a hybrid QEC-DD strategy.
-Figure 1: (a) Schematic of the experimental design: we encode a logical Bell state $\chi$ using the $[[4,2,2]]$ code and let it undergo either free evolution or LDD for a duration $\tau$.
+Definition 2 (SLDD).
+For any stabilizer group $\mathcal{S}$ and associated canonical group of logical operators $\mathcal{L}$, the stabilizer-logical dynamical decoupling (SLDD) group is
+$\tilde{\mathcal{G}}_{\text{SLDD}}=\pi(\mathcal{S}\mathcal{L}).$ (8)
+Note that $\mathcal{S}\mathcal{L}=\mathcal{N}(\mathcal{S})$ and thus independent of the choice of $\mathcal{L}$.
+The main weakness of SLDD is that it still incurs a cost that is exponential in the number of physical qubits $n$.
+This observation motivates us to consider a different hybrid approach that shifts more of the burden for handling correctable errors to the QEC code and uses only the logical operators of the code as DD pulses:
+Definition 3 (LDD).
+Pick a canonical group of logical operators $\mathcal{L}$. The logical dynamical decoupling (LDD) group is
+$\tilde{\mathcal{G}}_{\text{LDD}}\equiv\tilde{\mathcal{L}}=\pi(\mathcal{L}).$ (9)
+Compared to SLDD, this approach has the desired effect of reducing the DD group order to $T\leq 4^{k}$, independent of $n$.
+For the sake of completeness, we also mention the case of DD using only the stabilizers:
+Definition 4 (SDD).
+For any stabilizer group $\mathcal{S}$, the SDD group is
+$\tilde{\mathcal{G}}_{\text{SDD}}\equiv\tilde{\mathcal{S}}=\pi(\mathcal{S}).$ (10)
+SDD can be seen as an alternative to QEC since it decouples all detectable errors.
+However, it does not decouple any logical errors and is therefore not a particularly useful method when used in conjunction with error correction.
+It does play an important role in scenarios where QEC is not viable, such as in adiabatic quantum computing [79, 80].
+We can now classify the different decoupling groups by the errors they decouple.
+Theorem 1 (Decoupling by group type).
+  1. 1.
+    SDD decouples all detectable errors: $\langle{H_{SB}^{\mathcal{D}_{\star}}}\rangle_{\tilde{\mathcal{G}}_{\text{SDD}}}=0$.
+  2. 2.
+    LDD decouples all logical errors: $\langle{H_{SB}^{\mathcal{L}_{\star}}}\rangle_{\tilde{\mathcal{G}}_{\text{LDD}}}=0$.
+  3. 3.
+    SLDD decouples all errors: $\langle{H_{SB}^{\mathcal{L}_{\star}}+H_{SB}^{\mathcal{D}_{\star}}}\rangle_{\tilde{\mathcal{G}}_{\text{SLDD}}}=0$.
+Proof.
+It follows from 1 that it suffices to show that for each term in $H_{SB}^{\mathcal{S}_{\star}}$, $H_{SB}^{\mathcal{L}_{\star}}$, or $H_{SB}^{\mathcal{D}_{\star}}$, there exists $g\in\mathcal{G}$ it anticommutes with, where $\mathcal{G}$ is the corresponding decoupling group.
+Indeed:
+  1. 1.
+    SDD: Each detectable error anticommmutes with at least one element of $\mathcal{S}$.
+  2. 2.
+    LDD: Each nontrivial logical error anticommmutes with at least one element of $\mathcal{L}$.
+  3. 3.
+    Since $I\in\mathcal{L}$, the stabilizer $\mathcal{S}\subset\mathcal{G}_{\text{SLDD}}=\bigl{\{}\mathcal{S}L\mid L\in\mathcal{L}\bigr{\}}$.
+Thus, the group average of $H_{SB}^{\mathcal{L}_{\star}}+H_{SB}^{\mathcal{D}_{\star}}$ with respect to $\mathcal{G}_{\text{SLDD}}$ includes both $\mathcal{S}$, which decouples all detectable errors, and $\mathcal{L}$, which decouples all logical errors.
+Table 1 summarizes the relative resource cost of full Pauli group DD, SDD, LDD, and SLDD.
+Comparative DD resource cost
+Decoupling group $\mathcal{G}$ of pulses error type suppressed requires QEC
+$\tilde{\mathcal{P}}_{n}$ $2^{2n}$ all No
+SDD $2^{n-k}$ detectable No
+LDD $2^{2k}$ logical Yes
+SLDD $2^{n+k}$ detectable & logical No
+Table 1: The comparative resource cost for various DD methods in terms of the number of pulses required to suppress the most general system-bath Hamiltonian $H_{SB}$ [Eq.
+6] acting on $n$ qubits, and whether active error correction is required in addition.
+Full Pauli group DD and SLDD suppress all error mechanisms that can decohere code states, while LDD does not, as it leaves dealing with these error mechanisms for the QEC cycle.
+Consequently, LDD does not scale with $n$ but with $k$, whereas Pauli group DD and SLDD both scale with $n$.
+It is common to use only subgroups of the complete group of logical operators.
+In the context of standard (unencoded) DD, this is the case, e.g., when using single-axis sequences such as XX instead of the universal XY4 sequence.
+In the present context, we would like to consider subgroups of $\mathcal{N}(\mathcal{S})$ other than $\mathcal{S}$ and $\mathcal{L}$, which will similarly suppress selected subsets of errors.
+We refer to this as normalizer DD, since the corresponding DD group elements are generic normalizer elements, as opposed to canonical logical operators.
+Definition 5 (NDD).
+Normalizer DD (NDD) is any pulse sequence formed from using a subgroup of $\tilde{\mathcal{G}}_{\text{SLDD}}$ other than $\tilde{\mathcal{G}}_{\text{SDD}}$ or $\tilde{\mathcal{G}}_{\text{LDD}}$.
+NDD arises in our experiments, as described in Section III. We also discuss an example in the next subsection.
+It is important to emphasize that additional errors may and will generally be decoupled by each of the decoupling groups, beyond those mentioned in 1.
+For example, a detectable weight-$1$ Pauli operator that anticommutes with an element of $\mathcal{G}_{\text{LDD}}$ will be decoupled despite not being a logical error for any code with $d\geq 2$.
+Another important point is that since we are employing a Hamiltonian noise model, any terms that arise in second order perturbation theory, i.e., the $O(T^{2})$ terms we are neglecting in Eq.
+2, will mix $H_{SB}^{\mathcal{S}_{\star}}$, $H_{SB}^{\mathcal{L}_{\star}}$, and $H_{SB}^{\mathcal{D}_{\star}}$ through commutators.
+This means that the clean separation we have assumed between error types disappears at this higher order of perturbation theory.
+However, higher-order decoupling sequences are known that achieve suppression up to arbitrary order $q$ [i.e., leaving an $O(T^{q+1})$ unsuppressed error term in $U(T)$[72, 81, 82, 83, 84].
+For simplicity, in this work we restrict our attention to first-order sequences.
+II.4 Illustration of LDD and NDD: the $[[4,2,2]]$ code
+To illustrate the concepts discussed above, we analyze the example of the $[[4,2,2]]$ code, as it plays a key role in our experiments.
+The stabilizer group of the $[[4,2,2]]$ code is $\mathcal{S}=\langle XXXX,ZZZZ\rangle$.
+A canonical set of logical operators for the code can be chosen such that $\overline{X}_{1}=XIIX$, $\overline{X}_{2}=IIXX$, $\overline{Z}_{1}=IIZZ$ and $\overline{Z}_{2}=ZIIZ$, i.e., $\mathcal{L}=\langle i,XIIX,IIXX,IIZZ,ZIIZ\rangle$ (see Appendix B for additional pertinent details regarding this code).
+Thus, $\{XIIX,IIXX,IIZZ,ZIIZ\}$ is the generator set of the $16$-element LDD group $\tilde{\mathcal{G}}_{\text{LDD}}=\{\tilde{g}_{j}\}$.
+This LDD group (as well as any of the corresponding NDD groups) suppresses every logical error $E_{\text{log}}$ in the $602^{n-k}(4^{k}-1)=4\times 15=$ 60-element group $(\mathcal{N}(\mathcal{S})\setminus\mathcal{S})/\{\pm 1,\pm i\}$: the group-averaged logical system-bath Hamiltonian $\langle H_{SB}^{\mathcal{L}_{\star}}\rangle_{\mathcal{G}}$ [Eq.
+6c] vanishes, as every one of its terms is a logical operator that anticommutes with an element of $\mathcal{G}$.
+This leaves $1924^{n}-2^{n+k}=256-64=$ 192 detectable errors.
+These errors can be written compactly as $\mathcal{D}=\{(XIII\cdot\tilde{\mathcal{N}}(\mathcal{S}))\cup(YIII\cdot\tilde{\mathcal{N}}(\mathcal{S}))\cup(ZIII\cdot\tilde{\mathcal{N}}(\mathcal{S}))\}$, yielding $1923\times|\tilde{\mathcal{N}}(\mathcal{S})|=$ 192 terms.
+Each such error is detectable because it anticommutes with at least one element of $\mathcal{S}$.
+Of these, only the $12$ errors of the form $PQPP$, with $P\neq Q\in\{I,X,Y,Z\}$, are unsuppressed by LDD.
+Now suppose that instead of suppressing the full $H_{SB}^{\mathcal{L}_{\star}}$, the goal is to suppress just $H_{SB}^{\overline{Z}}\equiv\overline{Z}_{1}\otimes B_{1}^{\text{log}}+\overline{Z}_{2}\otimes B_{2}^{\text{log}}$.
+This scenario is particularly relevant to the experiments we describe below.
+In this case, it suffices to use the smaller LDD-subgroup $\langle\overline{X}_{1}\overline{X}_{2}\rangle$, since both $\overline{Z}_{1}$ and $\overline{Z}_{2}$ anticommute with $\overline{XX}\equiv\overline{X}_{1}\overline{X}_{2}$.
+The corresponding LDD cycle is then $\overline{XX}f_{\tau}\overline{XX}f_{\tau}$, where $f_{\tau}=\exp(-i\tau H_{SB}^{\overline{Z}})$.
+For crosstalk-related reasons that will become apparent below, it turns out that it is advantageous to replace the second $\overline{XX}$ by $S_{1}\overline{XX}=IXIX\equiv\overline{X^{\prime}X^{\prime}}$, where $S_{1}=XXXX\in\mathcal{S}$.
+This replacement yields the $4$-pulse NDD cycle $\overline{X^{\prime}X^{\prime}}f_{\tau}\overline{XX}f_{\tau}\overline{X^{\prime}X^{\prime}}f_{\tau}\overline{XX}f_{\tau}=(\overline{X^{\prime}X^{\prime}}f_{\tau}\overline{X^{\prime}X^{\prime}})(S_{1}f_{\tau}S_{1})(\overline{XX}f_{\tau}\overline{XX})(If_{\tau}I)$.
+This cycle is generated by the NDD group $\tilde{\mathcal{G}}=\langle\overline{XX},S_{1}\rangle=\{I,\overline{XX},S_{1},\overline{X^{\prime}X^{\prime}}\}$.
+Below, we refer to this decoupling group as NXX.
+As another example of NDD, consider the encoded XY4 sequence with the LDD group $\tilde{\mathcal{G}}=\langle\overline{XX},\overline{ZZ}\rangle$ and DD cycle $\overline{YY}f_{\tau}\overline{XX}f_{\tau}\overline{YY}f_{\tau}\overline{XX}f_{\tau}$.
+Again, for crosstalk-related reasons it will turn out to be advantageous to replace the canonical $\overline{YY}=YIYI$ by $\overline{Y^{\prime}Y^{\prime}}=S_{2}\overline{YY}$, where $S_{2}=YYYY\in\mathcal{S}$.
+The resulting DD cycle $\overline{Y^{\prime}Y^{\prime}}f_{\tau}\overline{XX}f_{\tau}\overline{Y^{\prime}Y^{\prime}}f_{\tau}\overline{XX}f_{\tau}$ is generated by the NDD group $\tilde{\mathcal{G}}=\langle\overline{XX},\overline{Z^{\prime}Z^{\prime}}\rangle=\{I,\overline{XX},\overline{Y^{\prime}Y^{\prime}},\overline{Z^{\prime}Z^{\prime}}\}$, where $\overline{Z^{\prime}Z^{\prime}}=S_{2}\overline{ZZ}$.
+Below, we refer to this decoupling group as NXY4.
+Figure 1: (a) Schematic of the experimental design: we encode a logical Bell state $\chi$ using the $[[4,2,2]]$ code and let it undergo either free evolution or NDD for a duration $\tau$.
 We then run the unencoding circuit for $\chi^{\prime}$ and measure the resulting bitstring.
@@
 The color code indicates that the encoded state has undergone either no error (black) or one of the logical errors (red, blue, green) indicated in (b).
-The DD sequences used in our experiments are (d) Robust Logical XX (RLXX), (e) Robust Logical XY4 (RLXY4), and (f) physically Staggered XY4 (SXY4).
-Each panel shows both the physical implementation (top) and the logical interpretation of each sequence.
-A tilde in (d) and (e) denotes a negative $\pi$ rotation, i.e., $X=R_{X}(\pi)=\exp(-i\pi\sigma^{x})$ $\tilde{X}=R_{X}(-\pi)=\exp(i\pi\sigma^{x})$, and likewise for $Y$.
-II The QEC-LDD Theorem
-The main weakness of SLDD is that it still incurs a cost that is exponential in the number of physical qubits $n$.
-Here, we present a different hybrid approach, shifting more of the burden for handling correctable errors to the QEC code and using only the logical operators as DD pulses $g_{j}$.
-We call this approach ‘logical DD’ (LDD).
-Compared to SLDD, this approach has the desired effect of reducing the single sequence time-cost to $T=4^{k}\tau$ for first-order decoupling, independent of $n$.
-Table 1 summarizes the relative resource cost of full Pauli group DD, SLDD, and LDD.
-Comparative DD resource cost
-Decoupling group $\mathcal{G}$ of pulses suppresses all errors afflicting code space / requires QEC
-Pauli group $\tilde{\mathcal{P}}_{n}$ $2^{2n}$ Yes / No
-SLDD $2^{n+k}$ Yes / No
-LDD $2^{2k}$ No / Yes
-Table 1: The comparative resource cost for various DD methods in terms of the number of pulses required to suppress the most general system-bath Hamiltonian $H_{SB}$ [Eq.
-5] acting on $n$ qubits, and whether active error correction is required in addition.
-Full Pauli group DD and SLDD suppress all error mechanisms which can decohere code states, while LDD does not, as it leaves dealing with these error mechanisms for the QEC cycle.
-Consequently, LDD does not scale with $n$ but with $k$, whereas Pauli group DD and SLDD both scale with $n$.
-II.1 Theorem statement and proof
-The LDD sequence is purely responsible for decoupling the logical errors acting within each syndrome space, but unlike SLDD, it no longer decouples some of the physical errors between the syndrome spaces.
-Correcting these errors is accomplished by syndrome measurements in conjunction with error recovery.
-For each code $\mathcal{C}$ there is a choice of which errors to correct: for each of the $12^{n-k}-$ 1 nontrivial syndromes, one may choose a recovery operation; this choice determines which errors are successfully corrected and which result in a logical error.
-Once such a choice is made, we call the errors that are successfully corrected “correctable”.
-We summarize this in the following theorem:
-Theorem 1 (QEC-LDD).
-Consider the most general system-bath Hamiltonian acting on $n$ qubits [Eq.
-5].
-Let $\mathcal{C}$ be an $[[n,k,d]]$ stabilizer code (with any $d\geq 1$).
-Let $\mathcal{L}=\langle i,\overline{X}_{j},\overline{Z}_{j}\rangle_{j=1}^{k}$ denote its group of canonical logical operators.
-Choose the decoupling group $\tilde{\mathcal{G}}=\mathcal{L}/\{\pm 1,\pm i\}$, henceforth the LDD group.
-Then $\tilde{\mathcal{G}}$ decouples all logical errors, that is, every term in $H_{SB}^{\mathcal{L}}$, as well as some errors in $H_{SB}^{\mathcal{E}}$.
-There is a choice of recovery operation such that all remaining non-decoupled errors in $H_{SB}^{\mathcal{E}}$ are correctable by $\mathcal{C}$.
-Conversely, for any recovery operation chosen in this way, only those errors that are not decoupled by $\tilde{\mathcal{G}}$ are correctable.
-Proof.
-We first establish when an error $E\in\mathcal{P}_{n}$ is decoupled.
-Let $\mathcal{H}=(\mathbb{C}^{2})^{\otimes n}$ be the Hilbert space of $n$ qubits.
-Then $\mathcal{C}\subset\mathcal{H}$ and all elements of $\mathcal{P}_{n}$ (and, hence, of $\mathcal{L}\subset\mathcal{P}_{n}$ and $\mathcal{S}\subset\mathcal{P}_{n}$) are unitary operators on $\mathcal{H}$.
-We let $\mathcal{G}=\mathcal{L}=\{\pm g_{j},\pm ig_{j}\}_{j=0}^{4^{k}-1}$.
-Consider a term in $H_{SB}$ corresponding to some $E\in\mathcal{P}_{n}$.
-It is decoupled if and only if there exists $g\in\mathcal{G}$ anticommuting with $E$.
-Indeed, then $g^{\dagger}Eg=-E$ and
-$\langle E\rangle_{\mathcal{G}}=\frac{1}{\left|\mathcal{G}\right|}\sum_{h\in\mathcal{G}}h^{\dagger}Eh=\frac{1}{\left|\mathcal{G}\right|}\sum_{h\in\mathcal{G}}(gh)^{\dagger}E{gh}=-\langle E\rangle_{\mathcal{G}},$ (6)
-so $\langle E\rangle_{\mathcal{G}}=0$.
-If there is no such $g$ then all elements of $\mathcal{G}$ commute with $E$ (because every two elements of $\mathcal{P}_{n}$ either commute or anticommute).
-Then all $g^{\dagger}Eg=E$ for all $g\in\mathcal{G}$, hence $\langle E\rangle_{\mathcal{G}}=E$ (i.e., the term is undecoupled).
-It follows that $\tilde{\mathcal{G}}$ decouples all logical errors, since every such error anticommutes with an element of $\mathcal{G}$.
-The errors in $H_{SB}^{\mathcal{E}}$ that anticommute with an element of $\mathcal{G}$ are also decoupled.
-All terms in $H_{SB}^{\mathcal{S}}$ are undecoupled since they all commute with $\mathcal{G}$.
-The key observation in the proof is that the statement of the theorem is invariant under Clifford maps, which we now show.
-Now consider any Clifford unitary $V\colon\mathcal{H}\to\mathcal{H}$.
-Let $\mathcal{C}^{\prime}=V\mathcal{C}$, $d^{\prime}\geq 1$ be the distance of the code $\mathcal{C}^{\prime}$, $\mathcal{G}^{\prime}=\mathcal{L}^{\prime}=V\mathcal{L}V^{\dagger}$, $\tilde{\mathcal{G}}^{\prime}=\mathcal{G}^{\prime}/\{\pm 1,\pm i\}$, $H_{SB}^{\prime}=VH_{SB}V^{\dagger}$, etc.
-For a recovery map $R_{s}$ (for every value $s$ of the syndrome) for $\mathcal{C}$ we can construct the corresponding recovery map $R_{s}^{\prime}=VR_{s}V^{\dagger}$ for $\mathcal{C}^{\prime}$.
-If the statement of the theorem is true for the original code $\mathcal{C}$, then it remains true for $\mathcal{C}^{\prime}$ because it only depended on the commutation relations between elements of $\mathcal{P}_{n}$ and those are preserved under the group homomorphism $P\mapsto VPV^{\dagger}$.
-By applying this observation to $V^{\dagger}$ instead of $V$ we observe that the converse also holds.
-In 1 in Appendix C, extending earlier methods for stabilizer code encoding [91, 92, 93], we show that for any code $\mathcal{C}$ and any choice of the canonical logical operators $\mathcal{L}$ there is a Clifford unitary $V$ (which we may call “unencoding circuit”) such that the corresponding code $\mathcal{C}^{\prime}$ consists of states of the form |$\ket{\psi}\otimes\ket{0}^{n-k}$, where |$\ket{\psi}\in(\mathbb{C}^{2})^{\otimes k}$, and $\mathcal{L}^{\prime}=V\mathcal{L}V^{\dagger}\simeq\mathcal{P}_{k}$ consists of Paulis $P\in\mathcal{P}_{n}$ acting trivially on the last $n-k$ qubits.
-We call such code $\mathcal{C}^{\prime}$ a trivial code: its logical qubits can be chosen to coincide with the first $k$ physical qubits and stabilizer generators can be chosen to be $Z_{k+1},\dots,Z_{n}$ enforcing that the last $n-k$ qubits are set to 00, and $d^{\prime}=1$.
-It follows that it is sufficient to prove the statement of the theorem for this trivial code.
-Hence, from now on we assume that $\mathcal{C}$ is the trivial code described above.
-$E\in\mathcal{P}_{n}$ is undecoupled if and only if it commutes with all Paulis acting on the first $k$ qubits.
-This is true if and only if $E$ acts trivially on the first $k$ qubits.
-(Equivalently, $E\in\mathcal{P}_{n}$ is decoupled if and only if acts non-trivially on one or more of the first $k$ qubits.) Measuring the stabilizers $Z_{k+1},\dots,Z_{n}$ is equivalent to measuring the last $n-k$ qubits.
-Any such measurement outcome $s=(s_{1},\dots,s_{n-k})$ corresponds to errors $E=SE_{s}$, where $E_{s}=\prod_{j=1}^{n-k}X_{k+j}^{s_{j}}$ and $S$ is an arbitrary stabilizer.
-A (unitary) recovery operation $R_{s}$ corrects such an error $E$ if and only if $R_{s}=S^{\prime}E_{s}$.
-In particular, if we apply $E_{s}$ corresponding to the measured syndrome $s$, we correct all non-decoupled errors, proving the first statement of the theorem.
-On the other hand, all such recovery maps act trivially on the first $k$ qubits, so if $E$ acts non-trivially on the first $k$ qubits (i.e.
-if $E$ is decoupled), it remains uncorrected by any such procedure.
-∎
-We emphasize a few key points:
-  • The claim holds even for $d=2$ (or $d=1$), i.e., an error detecting code becomes an error correcting code subject to LDD.
-  • In fact, the notion of correctable errors used here ignores the code distance altogether.
-For any distance, the QEC-LDD protocol has at least $3k$ non-correctable single-qubit errors (these are decoupled by LDD).
-[^1] Consider the group $\mathcal{L}$ and divide the physical qubits into three classes according to the action of the elements of $\mathcal{L}$ on a particular qubit.
-1.
-Every element of $\mathcal{L}$ acts as $I$ on that qubit.
-2.
-Every element of $\mathcal{L}$ acts as $I$ or one other Pauli ($X$, $Y$, or $Z$) on that qubit.
-3.
-For every Pauli in $\{I,X,Y,Z\}$ there is $P\in\mathcal{L}$ which acts like that Pauli on that qubit.
-Let $l_{1},l_{2},l_{3}$ be the number of physical qubits in each class.
-Restrict every element of $\mathcal{L}$ to the qubits from class 3.
-Such a restriction does not change the commutation relations.
-Let the resulting group be $\mathcal{L}^{\prime}$.
-We have $\mathcal{P}_{k}\simeq\mathcal{L}\simeq\mathcal{L}^{\prime}$, which in turn is embedded in $\mathcal{P}_{l_{3}}$.
-Therefore, $l_{3}\geq k$.
-The number of uncorrectable single-qubit errors is $2l_{2}+3l_{3}\geq l_{3}\geq 3k$.
-E.g., for the $[[4,2,2]]$ code $l_{2}=0,l_{3}=3,k=2$, $2l_{2}+3l_{3}=9>6=3k$.
-For example, in the standard distance-$d$ surface code there is a weight $1$ error, which can be “corrected” to become a weight $12d-$ 1 error.
-That does not contradict the theorem, because that error is suppressed by LDD.
-  • For each syndrome, the corresponding recovery operation is unique up to multiplication by stabilizers.
-  • All decoupled errors are not correctable; it is the role of LDD to suppress such errors.
-II.2 Illustration of the QEC-LDD Theorem: the $[[4,2,2]]$ code
-To illustrate the QEC-LDD construction, we analyze the example of the $[[4,2,2]]$ code (see Appendix A for a review of this code).
-The stabilizer group of this code is $\mathcal{S}=\langle XXXX,ZZZZ\rangle$.
-A set of logical operators for the code can be chosen such that $\overline{X}_{1}=XIIX$, $\overline{X}_{2}=IIXX$, $\overline{Z}_{1}=IIZZ$ and $\overline{Z}_{2}=ZIIZ$, i.e., $\mathcal{L}=\langle i,XIIX,IIXX,IIZZ,ZIIZ\rangle$.
-Up to a global phase, $\{XIIX,IIXX,IIZZ,ZIIZ\}$ is the generator set of the $16$-element decoupling group $\tilde{\mathcal{G}}=\{g_{j}\}$.
-This LDD group suppresses every logical error $E$ in the $4\times 16=64$-element $\mathcal{S}\mathcal{L}/\{\pm 1,\pm i\}$: the group-averaged logical system-bath Hamiltonian $\langle H_{SB}^{\mathcal{L}}\rangle_{\mathcal{G}}$ vanishes, as every one of its terms is a logical operator that anticommutes with an element of $\mathcal{G}$.
-The errors that are not suppressed are easily seen to be the $16$-element set =$\mathcal{E}=\{P\otimes Q\otimes P\otimes P:P,Q\in\{I,X,Y,Z\}\}$, since they commute with all elements of $\mathcal{G}$.
-Each such error is detectable because it anticommutes with at least one element of $\mathcal{S}$.
-Moreover, every such error is correctable since it is equal to the product of the stabilizer $P\otimes P\otimes P\otimes P$ and $I\otimes R\otimes I\otimes I$, where $R=Q^{\dagger}P$ can be determined (up to the global phase) from the value of the syndrome.
-The recovery operation is then the weight-one operator $I\otimes R\otimes I\otimes I$.
-Ignoring phases, the Pauli group on $4$ physical qubits, $\tilde{\mathcal{P}}_{4}$, has $2564^{4}=$ 256 elements.
-Of these, $64$ are the elements of $\mathcal{S}\mathcal{L}/\{\pm 1,\pm i\}$, $16$ are the unsuppressed errors $\mathcal{E}$, which leaves $176$ other elements $[P]=\{\pm P,\pm iP\}\in\tilde{\mathcal{P}}_{4}$ such that corresponding $P\in\mathcal{P}_{4}$ anticommutes with at least one element of $\mathcal{G}$ and is, therefore, suppressed by it.
-Appendix C illustrates the unencoding of the $[[4,2,2]]$ code.
-In this case, $\mathcal{P}_{2}$ decouples all errors on the first two (data) qubits and only $X_{3}$ or $X_{4}$ are left as undecoupled bit-flip errors on the last two qubits.
-II.3 Avoiding the use of virtual-$Z$ gates
-One of the limitations of our experimental setup is that physical $Z$ gates are unavailable: fixed-frequency transmons utilize virtual-$Z$ gates [94, 95].
-Such gates are problematic for DD [95].
-We now demonstrate that there are DD sequences satisfying the conditions of the QEC-LDD theorem which do not use any physical $Z$ pulses.
-To construct such a sequence, we first choose generators of $\mathcal{L}$ that do not involve physical $Z$ gates.
-An example choice would be $\mathcal{L}=\langle i,XIIX,IIXX,IIYY,YIIY\rangle$.
-One could follow the construction below with this choice.
-However, in order to improve the ability of the DD sequence to suppress $ZZ$ crosstalk we suggest a different choice.
-First, multiply the original generators of $\mathcal{L}$ (that is, $XIIX,IIXX,IIZZ,ZIIZ$) by the stabilizers $XXXX$, $XXXX$, $XXXX$, and $ZZZZ$, respectively, to obtain $\mathcal{L}^{\prime}=\langle i,IXXI,XXII,XXYY,IZZI\rangle$.
-Then, after some group operations, we can write the same group using a different set of generators containing no $Z$’s: $\mathcal{L}^{\prime}=\langle i,XXII,IIYY,IYIY,XIXI\rangle$.
-The DD sequence consisting of two repetitions of $XIXI$, $IYIY$, $XIXI$, $IIYY$, $XIXI$, $IYIY$, $XIXI$, $XXII$ then corresponds to the decoupling group /{±$\mathcal{L}^{\prime}/\{\pm 1,\pm i\}$, and involves no $Z$ gates.
-This sequence was obtained as follows: denote the generators of $\mathcal{L}$ (or $\mathcal{L}^{\prime}$) as $h_{1},h_{2},h_{3},h_{4}$ and choose any Gray code — a sequence of bitstrings $a_{0},\dots,a_{15}$ such that neighboring bitstrings differ in only a single digit.
-Finally, pick $g_{j}=h^{a_{j}}$ (for $j=0,\dots,15$), where $a_{j}$ is interpreted as a multiindex (i.e.
-$g_{j}=\prod_{k=1}^{4}h_{k}^{(a_{j})_{k}}$).
-The pulses of the DD sequence are $g_{j}g_{j+1}$ for $j=0,\dots,15$.
-However, in our experiments we found that a decoupling group generated by a subgroup of $\mathcal{L}$ and a subgroup of $\mathcal{S}$ works better in practice because of its robustness and ability to suppress crosstalk.
-The next section is dedicated to demonstrating experimentally how the hybrid QEC-LDD protocol results in a significant fidelity enhancement of entangled logical qubits.
-No DD XY4 UR6 UR8 UR10 UR18 RGA8a
-|$\ket{\Phi_{+}}$ $65.44\pm 0.24$ $83.94\pm 0.18$ $83.08\pm 0.19$ $83.11\pm 0.19$ $82.05\pm 0.19$ $81.43\pm 0.20$ $\boxed{84.20\pm 0.18}$
-|$\ket{\Phi_{-}}$ $61.23\pm 0.24$ $\boxed{83.78\pm 0.18}$ $81.92\pm 0.19$ $82.63\pm 0.19$ $81.07\pm 0.19$ $80.81\pm 0.18$ $83.34\pm 0.18$
-|$\ket{\Psi_{+}}$ $66.18\pm 0.23$ $\boxed{84.68\pm 0.18}$ $84.17\pm 0.18$ $83.37\pm 0.19$ $82.46\pm 0.19$ $81.21\pm 0.19$ $84.65\pm 0.18$
-|$\ket{\Psi_{-}}$ $65.43\pm 0.23$ $\boxed{86.57\pm 0.17}$ $85.20\pm 0.18$ $85.53\pm 0.18$ $84.17\pm 0.18$ $83.44\pm 0.17$ $86.38\pm 0.17$
-[…]
-III Experimental design and methodology
+The DD sequences used in our experiments are (d) Robust Normalizer XX (RNXX), (e) Robust Normalizer XY4 (RNXY4), and (f) physically Staggered XY4 (SXY4).
+Each panel shows both the physical implementation (top) and the logical interpretation of each sequence (bottom).
+A dagger in (d) and (e) denotes a negative $\pi$ rotation, i.e., $X=R_{x}(\pi)=\exp(-i\pi\sigma^{x}/2)$, $\tilde{X}=R_{x}(-\pi)=\exp(i\pi\sigma^{x}/2)$, and likewise for $Y$.
+In (d) the first column is the canonical $\overline{XX}=XIXI$ (see Appendix B), the second column is $\overline{X^{\prime}X^{\prime}}=S_{1}\overline{XX}$, where $S_{1}=XXXX\in\mathcal{S}$, etc.
+Likewise, in (e), the second column is $\overline{Y^{\prime}Y^{\prime}}=S_{2}\overline{YY}$, where the canonical $\overline{YY}=YIYI$ and $S_{2}=YYYY\in\mathcal{S}$.
+III Experimental design
+For our experiments, while sequences utilizing the full LDD/NDD group can be constructed (see Appendix C), we instead use a subgroup-NDD, generated by a subgroup of $\mathcal{L}$ and a subgroup of $\mathcal{S}$.
+This subgroup NDD works better in practice because of its robustness and ability to suppress crosstalk.
+This section is dedicated to demonstrating experimentally how our hybrid QED + subgroup-NDD protocol results in a significant fidelity enhancement of entangled logical qubits.
 The data for the experiments we report here were collected on two separate occasions from a total of $24$ sets of four-qubit experiments run on the ibm_kyiv [85] quantum processor.
@@
 Other errors of weight $\geq 2$, if present, would likewise correspond to logical errors.
-To demonstrate that LDD can suppress all logical errors, we design an experiment that enables the unequivocal detection of such errors.
+To demonstrate that NDD can suppress all logical errors, we design an experiment that enables the unequivocal detection of such errors.
 To this end, we use the two logical qubits of the $[[4,2,2]]$ code to create logical Bell states |$\ket{\Phi_{\pm}}=(\ket{\overline{00}}\pm\ket{\overline{11}})/\sqrt{2}$ and |$\ket{\Psi_{\pm}}=(\ket{\overline{01}}\pm\ket{\overline{10}})/\sqrt{2}$.
-The corresponding encoding circuits $U_{\text{enc},\chi}$, where $\chi\in\{\ket{\Phi_{\pm}},\ket{\Psi_{\pm}}\}$, create two copies of physical Bell states, followed by the application of a logical controlled-NOT ($\overline{\text{CNOT}}$) to generate the logical Bell states (see Appendix A).
+The corresponding encoding circuits $U_{\text{enc},\chi}$, where $\chi\in\{\ket{\Phi_{\pm}},\ket{\Psi_{\pm}}\}$, create two copies of physical Bell states, followed by the application of a logical controlled-NOT ($\overline{\text{CNOT}}$) to generate the logical Bell states (see Appendix B).
+In principle, it would be possible to detect specific logical errors using other logical states (e.g., |$\ket{\overline{00}}$ would allow us to detect $X$- or $Y$-type logical errors), but we specifically opt for logical Bell states as they have shorter encoding circuits.
+Additionally, the four logical Bell states form an orthonormal basis and the logical errors we are interested in permute this basis.
 To estimate the fidelity of a logical Bell state $\chi$, we start from the physical ground state |$\ket{0000}$, encode into the state of interest by applying $U_{\text{enc},\chi}$, and then let it evolve for some time $\tau$, either freely or subject to DD.
-If after time $\tau$ we unencode by applying $U_{\text{enc},\chi}^{\dagger}$, then measuring the bitstring $0000$ would indicate that no error occurred.
+After time $\tau$ we unencode by applying $U_{\text{enc},\chi}^{\dagger}$, and measure all qubits.
+Measuring the bitstring $0000$ would indicate that no error occurred.
 If, instead, we obtain any of the odd-Hamming weight bitstrings, this would signal detection of a physical error, as such states are not in the codespace.
@@
 This procedure can be used to estimate the probability of finding the ground state as the empirical fraction of ground state measurement outcomes, which is also the fidelity of the logical Bell state $\chi$.
-Figure 2: Circuit schematic for encoding |$\ket{\Phi_{+}}$ and unencoding |$\ket{\Phi_{-}}$ with physical and logical DD.
-Physical XY4 dynamical decoupling sequences (yellow boxes) are inserted into the idle gaps of the circuit.
-The code subspace is protected with LDD sequences.
-This is the DD-protected version of the circuit shown in Fig.
-1(a).
-Figure 3: Performance of LDD (dataset 1).
-The experiments encode the logical Bell states |$\ket{\Phi_{+}}$ in (a) and (b), and |$\ket{\Psi_{+}}$ in (c) and (d).
-We then unencode in the logical Bell state related through $\overline{Z}_{i}$ on either of the logical qubits; i.e., |$\ket{\Phi_{+}}\mapsto\ket{\Phi_{-}}$ in (a) and (b), and |$\ket{\Psi_{+}}\mapsto\ket{\Psi_{-}}$ in (c) and (d).
-In this setting, the frequency of $0110$ outcomes [(a) and (c)] is a measure of the fidelity of the originally encoded state, while $0000$ corresponds to the detection of a $\overline{Z}_{i}$ error [(b) and (d)].
-Without any DD, the fidelity is low (green in (a) and (c)) and the probability of logical errors is high [green in (b) and (d)].
-Standard LDD sequences (LXX and LXY4; dashed light red and blue) improve the fidelity and logical error probability at short times ($\lesssim 15\mu$s) but their performance declines at longer times: they exhibit coherent errors as seen in the oscillations of the corresponding curves.
-Robust sequences (RLXX and RLXY4; solid dark red and blue) are the top performers; their ability to suppress the $\overline{Z}_{i}$ error is particularly noteworthy.
-Error bars are $1\sigma$ standard deviation after bootstrapping the data.
-To obtain the fidelity of a logical error, we first observe that each of the logical Bell states is related to the other three logical Bell states through a specific logical error: $\chi^{\prime}=\overline{O}\chi$, where $\overline{O}\in\{I,\overline{X}_{i},\overline{Y}_{i},\overline{Z}_{i}\}$ and $i=1$ or $2$; i.e., a logical operator that acts on either of the two logical qubits.
+No DD XY4 UR6 UR8 UR10 UR18 RGA8a
+|$\ket{\Phi_{+}}$ $65.44\pm 0.24$ $83.94\pm 0.18$ $83.08\pm 0.19$ $83.11\pm 0.19$ $82.05\pm 0.19$ $81.43\pm 0.20$ $\boxed{84.20\pm 0.18}$
+|$\ket{\Phi_{-}}$ $61.23\pm 0.24$ $\boxed{83.78\pm 0.18}$ $81.92\pm 0.19$ $82.63\pm 0.19$ $81.07\pm 0.19$ $80.81\pm 0.18$ $83.34\pm 0.18$
+|$\ket{\Psi_{+}}$ $66.18\pm 0.23$ $\boxed{84.68\pm 0.18}$ $84.17\pm 0.18$ $83.37\pm 0.19$ $82.46\pm 0.19$ $81.21\pm 0.19$ $84.65\pm 0.18$
+|$\ket{\Psi_{-}}$ $65.43\pm 0.23$ $\boxed{86.57\pm 0.17}$ $85.20\pm 0.18$ $85.53\pm 0.18$ $84.17\pm 0.18$ $83.44\pm 0.17$ $86.38\pm 0.17$
+[…]
+To obtain the fidelity of a logical error, we first observe that each of the logical Bell states is related to the other three logical Bell states through a specific logical error: $\chi^{\prime}=\overline{O}\chi$, for $\overline{O}\in\{\overline{X}_{i},\overline{Y}_{i},\overline{Z}_{i}\}$ and $i=1$ or $2$; i.e., a logical operator that acts on either of the two logical qubits.
+More generally, any other combinations of logical weight-$1$ errors also map the input Bell state to the other Bell states.
+That is, for $\bar{p},\bar{q}\in\{\overline{X},\overline{Y},\overline{Z}\}$, we have $\chi^{\prime}=(\bar{p}\otimes\bar{q})\chi=I\otimes(\bar{p}\bar{q})\chi=(\bar{p}\bar{q})\otimes I\chi$, up to overall phases.
+We note that the following two-qubit logical errors leave the logical Bell states unchanged and therefore do not influence the experimental outcome: $\overline{O}\chi=\chi$, for $\overline{O}\in\{II,\overline{X}\overline{X},\overline{Y}\overline{Y},\overline{Z}\overline{Z}\}$.
 We thus proceed as follows: instead of applying $U_{\rm{enc},\chi}^{\dagger}$, at $t=\tau$ we deliberately unencode into a different logical Bell state $\chi^{\prime}$ [see Fig.
@@
 From here on, we use the notation $\chi\mapsto\chi^{\prime}$ to denote the procedure of preparing the encoded logical Bell state $\chi$ and unencoding into $\chi^{\prime}$, i.e., of using the encoding unitary $U_{\text{enc},\chi}$ and the unencoding unitary $U^{\dagger}_{\text{enc},\chi^{\prime}}$.
-We note that logical state tomography is an alternative method for learning about the performance of LDD, and in particular, for certifying the entanglement of logical Bell states.
+We note that logical state tomography is an alternative method for learning about the performance of LDD or NDD, and in particular, for certifying the entanglement of logical Bell states.
 However, the $[[4,2,2]]$ code imposes some additional challenges in performing logical tomography, which precluded its use in our work; see Appendix D for details.
-III.1 Physical dynamical decoupling improves logical Bell state fidelity
-We first show that we can substantially improve the logical Bell fidelity by padding the idle gaps of the encoding circuits with physical (as opposed to logical) DD sequences.
+Figure 2: Circuit schematic for encoding |$\ket{\Phi_{+}}$ and unencoding |$\ket{\Phi_{-}}$ with physical and logical DD.
+Physical XY4 dynamical decoupling sequences (yellow boxes) are inserted into the idle gaps of the circuit.
+The code subspace is protected with NDD sequences.
+This is the DD-protected version of the circuit shown in Fig.
+1(a).
+Figure 3: Performance of NDD (dataset 1).
+The experiments encode the logical Bell states |$\ket{\Phi_{+}}$ in (a) and (b), and |$\ket{\Psi_{+}}$ in (c) and (d).
+We then unencode in the logical Bell state related through $\overline{Z}_{i}$ on either of the logical qubits; i.e., |$\ket{\Phi_{+}}\mapsto\ket{\Phi_{-}}$ in (a) and (b), and |$\ket{\Psi_{+}}\mapsto\ket{\Psi_{-}}$ in (c) and (d).
+In this setting, the frequency of $0110$ outcomes [(a) and (c)] is a measure of the fidelity of the originally encoded state, while $0000$ corresponds to the detection of a $\overline{Z}_{i}$ error [(b) and (d)].
+Without any DD, the fidelity is low [green in (a) and (c)] and the probability of logical errors is high [green in (b) and (d)].
+Standard NDD sequences (NXX and NXY4; dashed light red and blue) improve the fidelity and logical error probability at short times ($\lesssim 15\mu$s) but their performance declines at longer times: they exhibit coherent errors as seen in the oscillations of the corresponding curves.
+Robust sequences (RNXX and RNXY4; solid dark red and blue) are the top performers; their ability to suppress the $\overline{Z}_{i}$ error is particularly noteworthy.
+Error bars are $1\sigma$ standard deviation after bootstrapping the data.
+IV Experimental Results
+IV.1 Physical dynamical decoupling improves logical Bell state fidelity
+We first show that we can substantially improve the logical Bell fidelity by padding the idle gaps of the encoding circuits with physical (as opposed to logical or normalizer) DD sequences.
 An idle gap is a temporal circuit segment during which no gates are applied.
 Such gaps occur, e.g., when a pair of qubits is involved in a two-qubit gate that takes much longer than a single-qubit gate simultaneously being applied to another qubit; the latter is then idle after the completion of the single-qubit gate, while awaiting the completion of the two-qubit gate.
-As shown in Fig.
-2, we inserted various DD sequences into the idle gaps of Bell state circuits.
-This includes XY4 [40], universally robust sequences URn [62] for $n=6,8,10,18$, and RGA8a [61]; the former are illustrated along with robust variants in Fig.
-1(d-f).
+As shown in Fig. 2, we inserted various DD sequences into the idle gaps of Bell state circuits. This includes XY4 [40], universally robust sequences URn [62] for $n=6,8,10,18$, and RGA8a [61].
+Since the URn sequences rely heavily on the use of both positive and negative rotations, it is important to clarify that henceforth, we use the notation $X\equiv R_{x}(\pi)=\exp(-i\pi\sigma^{x}/2)$, $\tilde{X}\equiv R_{x}(-\pi)=\exp(i\pi\sigma^{x}/2)$, and likewise for $Y$, to denote physical DD pulses.
+That is, unlike Section II, the symbols $X$ and $Y$ no longer denote the Pauli $\sigma^{x}$ and $\sigma^{y}$ operators.
+The reason this matters is that physical DD pulses suffer from control errors (both axis-angle and angle-magnitude errors), as well as errors due to the presence of the system-bath interaction while the pulse is on.
+In other words, the notation $R_{x}(\pi)$, etc., hides the fact that in reality the exact rotation axis deviates from $x$ and the exact rotation angle deviates from $\pi$.
+When these deviations are accounted for, DD sequences that are formally identical due to phase cancellation, such as $XX$ and $X\tilde{X}$, result in different operations at the physical level.
+For more details, see Appendix E.
 We performed the experiments (during the week of August 12th 2024) on fourteen different sets of qubits of ibm_kyiv and report the average fidelity values in Table 2.
@@
 In all, these results demonstrate that logical Bell state preservation can benefit significantly from physical DD.
-Having established the utility of physical DD, we proceed to combine it with LDD in the next section.
+Having established the utility of physical DD, we proceed to combine it with NDD in the next section.
 Figure 4: Detection probability (i.e., probability of the $0000$ bitstring) of logical $X$, $Y$, and $Z$-type errors (dataset 1).
@@
 These experiments are performed without DD.
-III.2 Logical error suppression and detection by LDD
-For all our experimental results, ‘No DD’ refers to $[[4,2,2]]$ encoding without any physical or logical DD.
-In all other experiments, we use physical XY4 to pad all the encoding and unencoding circuits gaps, and combine them with various flavors of LDD.
+IV.2 Logical error suppression and detection by NDD
+For all our experimental results, “No DD” refers to $[[4,2,2]]$ encoding without any physical or logical DD.
+In all other experiments, we use physical XY4 to pad all the encoding and unencoding idle gaps, and combine them with various flavors of NDD.
 This choice allows us to clearly assess the improvements introduced by DD, compared to relying solely on the code’s error detection capabilities.
-III.2.1 Logical $Z$ errors
+IV.2.1 Logical $Z$ errors
 We start by gradually increasing the time delay $\tau$ (up to 55 $\mu$s) between the encoding and unencoding without DD.
 This situation is relevant in the context of QEC experiments.
-For example, one could prepare an encoded qubit and then leave it idle while other logical operations are applied to other encoded qubits [19].
+For example, one could prepare an encoded qubit and then leave it to idle while other logical operations are applied to other encoded qubits [19].
 We perform the $\overline{Z}_{i}$ error detection using the logical Bell states |$\ket{\Psi_{+}}$ and |$\ket{\Phi_{+}}$ as discussed in Section III, and display the results in Fig.
@@
 As can be seen in Fig.
-3(a) and (c), without LDD the free evolution fidelity (denoted as No DD) decays rapidly and exhibits $ZZ$ crosstalk-induced oscillations.
-Fig.
-3(b) and (d) show that logical $Z$ errors accumulate over time.
-We next use two types of LDD sequences to suppress this effect: Logical XX (LXX) and Logical XY4 (LXY4).
-Physical-level schematics of these LDD sequences are shown in the boxed four-pulse sequences of Fig.
+3(a) and (c), without NDD the free evolution fidelity (denoted as No DD) decays rapidly and exhibits $ZZ$ crosstalk-induced oscillations.
+Figure 3(b) and (d) show that logical $Z$ errors accumulate over time.
+We next use two types of NDD sequences to suppress this effect: Normalizer XX (NXX) and Normalizer XY4 (NXY4).
+Physical-level schematics of these NDD sequences are shown in the boxed four-pulse sequences of Fig.
 1(d) and (e), respectively (disregarding the tilde notation).
@@
 1.
-We generate both LXX and LXY4 using the native logical operations of the $[[4,2,2]]$ code.
+We generate both NXX and NXY4 using the native logical operations of the $[[4,2,2]]$ code.
 For example, in the upper part of Fig.
-1(d), reading the first column from top to bottom yields $XIXI=\overline{XX}$, which is the $\bar{X}\bar{X}$ column at the bottom of Fig.
+1(d), reading the first (second) column from top to bottom yields $XIXI=\overline{XX}$ ($\overline{X^{\prime}X^{\prime}}=IXIX=S_{1}\overline{XX}$; $S_{1}=XXXX$), which is the $\bar{X}\bar{X}$ ($\bar{X}^{\prime}\bar{X}^{\prime}$) column at the bottom of Fig.
 1(d).
-Reading the pulse sequences from left to right, the staggering (appearance of delays as indicated by the identity operations) is deliberately introduced to suppress crosstalk at the physical level [65]; see Appendix B for more details.
-The corresponding results are shown as dashed lines in Fig.
+Reading the pulse sequences from left to right, the staggering (appearance of delays as indicated by the identity operations) is deliberately introduced to suppress crosstalk at the physical level [65]; see Appendix E for more details.
+The corresponding results are denoted as NXX (red) and NXY4 (dark green) in Fig.
 3.
-At short times ($\lesssim 15\mu$s), these results are better than those without DD because some errors (including $ZZ$ crosstalk) are suppressed by both LXX and LXY4.
+NXX is a single-axis non-universal sequence, while NXY4 is “nearly” universal: it suppresses all logical errors except its own generating set $\{XIXI,IXIX,YIYI,IYIY\}$.
+At short times ($\lesssim 15\mu$s), these results are better than those without DD because some errors (including $ZZ$ crosstalk) are suppressed by both NXX and NXY4.
 However, at longer times, the benefit is lost and, moreover, small oscillations appear that indicate the presence of coherent errors  [89].
-To overcome this, we create robust versions of LXY4 and LXX by ensuring that all physical qubits in the code undergo physical DD sequences robust to pulse errors.
+To overcome this, we create robust versions of NXY4 and NXX by ensuring that all physical qubits in the code undergo physical DD sequences robust to pulse errors.
 Specifically, we use the universally robust (UR) sequence family [62], and ensure that each physical qubit undergoes a UR4 sequence, i.e., $X\tilde{X}\tilde{X}X$ or $Y\tilde{Y}\tilde{Y}Y$, where a tilde denotes an $X$ or $Y$ rotation by $-\pi$ instead of $\pi$.
-These robust versions, which we call RLXX and RLXY4, are the full sequences shown in Fig.
+These robust versions, which we call RNXX and RNXY4, are the full sequences shown in Fig.
 1(d) and (e).
-The performance of these sequences is shown by the solid lines in Fig.
-3 and exhibits a significant improvement.
-Notably, the logical Bell state fidelities decay more slowly and without oscillations (panels (a) and (c)), and the logical $Z$ errors are strongly suppressed (panels (b) and (d)).
+The performance of RNXX (yellow) and RNXY4 (blue), as seen in Fig.
+3, exhibits a significant improvement.
+Notably, the logical Bell state fidelities decay more slowly and without oscillations [panels (a) and (c)], and the logical $Z$ errors are strongly suppressed [panels (b) and (d)].
 In addition, we apply the physical staggered XY4 (SXY4) sequence.
@@
 This sequence, shown in Fig.
-1(f), also corresponds to an LDD sequence for the $[[4,2,2]]$ code (bottom of panel (f)).
-However, it is not a universal logical sequence, but rather a purely logical-$Z$-error suppressing sequence.
+1(f), also corresponds to a non-universal purely logical-$Z$-error suppressing, NDD sequence for the $[[4,2,2]]$ code [bottom of panel (f)].
 Interestingly, Fig.
-3 shows that SXY4 performs on par with RLXX and RLXY4.
+3 shows that SXY4 performs on par with RNXX and RNXY4.
 This finding signals that the dominant logical errors are of $Z$-type.
@@
 […]
-III.2.2 Logical $X$ and $Y$ errors
+IV.2.2 Logical $X$ and $Y$ errors
 So far, we have only discussed the detection and suppression of $Z$-type logical errors.
@@
 This confirms that $ZZ$ crosstalk is the main source of logical errors.
-III.2.3 LDD with postselection
-Having implemented LDD, we can further improve the results by using the error detection capability of the $[[4,2,2]]$ code, which allows us to perform postselection.
+IV.2.3 NDD with postselection
+Having implemented NDD, we can further improve the results by using the error detection capability of the $[[4,2,2]]$ code, which allows us to perform postselection.
 We do so by discarding any measurement outcome outside the logical basis (which would be the result of a physical error); i.e., we only keep measurement results corresponding to the bitstrings $\{0000,0110,1001,1111\}$.
+Doing so leads to our central experimental result, the demonstration of high-fidelity entangled logical qubits, shown in Fig.
+5.
 […]
-III.2.4 Beyond breakeven and the state-of-the-art
+IV.2.4 Beyond breakeven and the state-of-the-art
 Fig.
@@
 The former (‘Best Physical No DD’) exhibits strong crosstalk-induced oscillations with an overall fidelity comparable to that of the mean logical encoded Bell pairs case without DD (No DD+PS).
-Adding SXY4 significantly improves the fidelity and outperforms even the mean fidelity of logical Bell pairs with LDD (RLXY4).
-This shows that LDD by itself is not better than working with physical qubits and a crosstalk-robust DD sequence.
-However, physical Bell pairs with SXY4 are significantly worse than RLXY4+PS, i.e., the case of LDD with postselection on the results of the $[[4,2,2]]$ code.
-This constitutes clear evidence of beyond-breakeven performance for our QEC-LDD strategy.
-[…]
+Adding SXY4 significantly improves the fidelity and outperforms even the mean fidelity of logical Bell pairs with NDD (RNXY4).
+This shows that NDD by itself is not better than working with physical qubits and a crosstalk-robust DD sequence.
+However, physical Bell pairs with SXY4 are significantly worse than RNXY4+PS, i.e., the case of NDD with postselection on the results of the $[[4,2,2]]$ code.
+This constitutes clear evidence of beyond-breakeven performance for our QEC-NDD strategy.
 […]
@@
 Figure 6: The percentage of data discarded per circuit as a function of the delay time between encoding and unencoding (dataset 2).
-While LDD suppresses logical errors, it also reduces certain physical errors, as evidenced by the lower data discard rate when using LDD (i.e., fewer physical errors are detected).
-III.3 Physical error suppression and detection by LDD
-The LDD sequences suppress not only logical errors but physical errors as well.
-To see this, consider, e.g., the LXX sequence [Fig.
+While NDD suppresses logical errors, it also reduces certain physical errors, as evidenced by the lower data discard rate when using NDD (i.e., fewer physical errors are detected).
+IV.3 Physical error suppression and detection by NDD
+As explained in the discussion following 1 and in Section II.4, the NDD sequences suppress not only logical errors but physical errors as well.
+To see this, consider, e.g., the RNXX sequence [Fig.
 1(d)].
-As described in Section III.2.1, reading the pulse sequence vertically, each time step of this sequence operates in the logical subspace as $\overline{XX}$.
-Simultaneously, reading the sequence horizontally over a complete round of LDD, all four physical qubits undergo the physical XX sequence (i.e., $X-\tau-X-\tau$), which suppresses the set of physical errors $\{Y_{i},Z_{i}\}_{i=1}^{4}$.
-Similarly, for the RLXY4 sequence [Fig.
+As described in Section IV.2.1, reading the pulse sequence vertically, each time step of this sequence operates in the logical subspace as $\overline{XX}$.
+Simultaneously, reading the sequence horizontally over a complete round of NDD, all four physical qubits undergo the physical XX sequence (i.e., $X-\tau-X-\tau$), which suppresses the set of physical errors $\{Y_{i},Z_{i}\}_{i=1}^{4}$.
+Similarly, for the RNXY4 sequence [Fig.
 1(e)], each physical qubit undergoes UR4.
@@
 One way to gauge the impact of this suppression of physical errors is shown in Fig.
-6, which displays the percentage of discarded data per circuit as a result of postselection over $55\mu$s (out of $4000$ shots), comparing No DD to RLXY4 for the two logical Bell states |$\ket{\Phi_{+}}$ and |$\ket{\Psi_{+}}$.
+6, which displays the percentage of discarded data per circuit as a result of postselection over $55\mu$s (out of $4000$ shots), comparing No DD to RNXY4 for the two logical Bell states |$\ket{\Phi_{+}}$ and |$\ket{\Psi_{+}}$.
 The results are obtained through bootstrapping across the ten sets of four qubits used in these experiments.
-Significantly less data are discarded with RLXY4 than without DD.
-Since discarded data correspond to the detection of physical errors, this means that LDD not only enhances the fidelity of logical Bell states by suppressing logical errors, but also reduces the occurrence of physical errors.
+Significantly less data is discarded with RNXY4 than without DD.
+Since discarded data corresponds to the detection of physical errors, this means that NDD not only enhances the fidelity of logical Bell states by suppressing logical errors, but also reduces the occurrence of physical errors.
 We can go further and use Algorithmic Error Tomography [54] to identify specific physical error types.
@@
 7 we show the corresponding relative bitstring counts, which are measures of detecting either of these errors.
-Additionally, we show the $0000$ and $0110$ relative bitstring counts, which correspond to the logical fidelity and the logical $Z$ error, respectively.
-Without LDD (dashed), we observe that the logical fidelity is low and both logical and physical errors increase and oscillate.
-With DD (solid), the RLXY4 sequence strongly suppresses both types of errors.
-These results demonstrate that, as claimed, LDD sequences suppress both physical and logical errors.
+Additionally, we show the $0110$ and $0000$ relative bitstring counts, which correspond to the logical fidelity and the logical $Z$ error, respectively.
+Without DD (dashed), we observe that the logical fidelity is low, and both logical and physical errors increase and oscillate.
+With DD (solid), the RNXY4 sequence strongly suppresses both types of errors.
+These results demonstrate that, as claimed, NDD sequences suppress both physical and logical errors.
 Figure 7: The relative count of the $\{0000,0110,0100,0010\}$ bitstrings in the |$\ket{\Phi_{+}}\mapsto\ket{\Phi_{-}}$ experiments (dataset 2).
-The two even-weight bitstrings correspond to the logical fidelity ($0000$) and logical $Z$ error ($0110$), and the two odd-weight bitstrings correspond to the detection of physical $\{Z_{i}\}_{i=1}^{4}$ errors.
-Dashed (solid) lines correspond to no DD (RLXY4).
-The LDD sequence, in addition to suppressing the logical errors and improving the logical fidelity, also strongly suppresses the occurrence of these physical errors.
-IV Discussion
+The two even-weight bitstrings correspond to the logical fidelity ($0110$) and logical $Z$ error ($0000$), and the two odd-weight bitstrings correspond to the detection of physical $\{Z_{i}\}_{i=1}^{4}$ errors.
+Dashed (solid) lines correspond to no DD (RNXY4).
+The NDD sequence, in addition to suppressing the logical errors and improving the logical fidelity, also strongly suppresses the occurrence of these physical errors.
+V Discussion
 The operation of QEC codes is adversely affected by the occurrence of logical errors that the code cannot detect or correct.
-Here, we have shown how to combine QEC with dynamical decoupling implemented in terms of the logical operators of the code, resulting in a hybrid QEC-LDD strategy that is significantly more effective than either QEC or LDD alone.
-We designed our LDD sequences to simultaneously perform logical error suppression and to be robust DD sequences at the physical level, resistant to both control errors and crosstalk.
-Our results, using the $[[4,2,2]]$ code and IBM transmon qubits, demonstrate a beyond-breakeven fidelity of entangled logical qubits.
+Here, we have shown how to combine QEC with dynamical decoupling implemented in terms of the normalizer elements of the code, resulting in a hybrid QEC-NDD strategy that is significantly more effective than either QEC or NDD alone.
+We designed our NDD sequences to simultaneously perform logical error suppression and to be robust DD sequences at the physical level, resistant to both control errors and crosstalk.
+As an added benefit, these sequences suppress many physical errors as well.
+Our results using the $[[4,2,2]]$ code and IBM transmon qubits, featured in Fig.
+5, demonstrate a beyond-breakeven fidelity of entangled logical qubits.
+The fidelities we report are the highest to date for entangled logical qubits using superconducting qubits.
 Our findings address a need along the path toward fault-tolerant quantum computation: keeping codes relatively small and nimble while still effectively handling logical errors.
-Future research should aim to optimize LDD sequences tailored to specific codes and integrate QEC-LDD into quantum algorithms.
-Another interesting future direction is optimization of QEC-LDD for tunable-coupler transmon devices; we present preliminary results in Appendix E.
+Future research should aim to optimize NDD sequences tailored to specific codes and integrate QEC-NDD into quantum algorithms.
+Another interesting future direction is the optimization of QEC-NDD for tunable-coupler transmon devices; we present preliminary results in Appendix F.
 Acknowledgements.
@@
 Government.
-This research was conducted using IBM Quantum Systems provided through University of Southern California’s IBM Quantum Innovation Center.
+This research was conducted using IBM Quantum Systems provided through the University of Southern California’s IBM Quantum Innovation Center.
 The views expressed are those of the authors and do not reflect the official policy or position of IBM or the IBM Quantum team.
-Appendix A [[4,2,2]] code
+Appendix A Quantum codes and logical errors
+A general $[[n,k,d]]$ code encodes $k$ logical qubits into $n$ physical qubits with distance $d$ [92].
+The weight $w$ of an error is the number of physical qubits it affects simultaneously.
+An $[[n,k,d]]$ stabilizer code corrects every error whose weight satisfies $w\leq t=\bigl{\lfloor}\tfrac{d-1}{2}\bigr{\rfloor}$ and detects every error whose weight $w\leq d-1$.
+A $d=2$ code is a pure quantum error detection (QED) code, and a distance-$d$ QEC code can always be used as a QED code for errors of weight $<d$.
+The code distance is the minimum number of physical qubits that must experience an error to cause an undetectable logical error, i.e., an error forming a logical operation inside the code space.
+These logical errors can either be inherently present or result from the accumulation of lower-weight errors over time.
+For example, in the context of superconducting qubits with fixed-frequency couplers, a prevalent challenge is $ZZ$ crosstalk [93, 94], which inherently introduces weight-two errors that can present as logical errors for distance-$2$ codes.
+[…]
+For a code $\mathcal{C}\subset(\mathbb{C}^{2})^{\otimes n}$, the stabilizer group $\mathcal{S}$ is uniquely defined as
+$\mathcal{S}=\{P\in\mathcal{P}_{n}:\forall\ket{\psi}\in\mathcal{C},P\ket{\psi}=\ket{\psi}\}.$ (11)
+The normalizer (or centralizer) of $\mathcal{S}$ in $\mathcal{P}_{n}$ is the group $\mathcal{N}(\mathcal{S})$ of Pauli operators commuting with all elements of $\mathcal{S}$ [^2] The normalizer $\mathcal{N}(\mathcal{S})=\mathcal{N}_{\mathcal{G}}(\mathcal{S})$ of $\mathcal{S}\subseteq\mathcal{G}$ is defined as $\{g\in\mathcal{G}\text{ s.t.
+}g\mathcal{S}g^{-1}=\mathcal{S}\}$, and the centralizer $\mathcal{Z}(\mathcal{S})=\mathcal{Z}_{\mathcal{G}}(\mathcal{S})$ of $\mathcal{S}\subseteq\mathcal{G}$ is defined as $\{g\in\mathcal{G}\text{ s.t.
+}\forall s\in\mathcal{S},gs=sg\}$.
+These two notions coincide in the case of stabilizer groups.
+The normalizer is also uniquely defined.
+However, the group of canonical logical operators is not unique: one can choose arbitrary stabilizers and multiply the canonical generators $\bar{X}_{j},\bar{Z}_{j}$ by those to obtain other choices of the group $\mathcal{L}$.
+Another important set is $\mathcal{D}=\mathcal{P}_{n}\setminus\mathcal{N}(\mathcal{S})$, the set of detectable errors.
+Since every pair of elements of $\mathcal{P}_{n}$ either commute or anticommute, and $\mathcal{N}(\mathcal{S})$ contains all elements of $\mathcal{P}_{n}$ that commute with $\mathcal{S}$, every element of $\mathcal{D}$ must anticommute with at least one element of $\mathcal{S}$.
+Appendix B [[4,2,2]] code
 The $[[4,2,2]]$ code is an error detection code that encodes $k=2$ logical qubits into $n=4$ physical qubits [95].
@@
 Using the same definitions, we have $\overline{\text{CNOT}}_{12}=\text{SWAP}_{12}$ and similarly, $\overline{\text{CNOT}}_{21}=\text{SWAP}_{23}$, i.e., we can perform logical CNOTs by swapping the physical qubits.
-Our particular choice of logical operators is motivated by the fact that $\overline{XX}$ and $\overline{YY}$ (i.e., the logical operators we use to implement the LDD sequences) have a natural staggering of their physical $X$ and $Y$ gates.
-In other words, in the implementation of the logical operators comprising LDD, nearest-neighbor qubits are always interleaved with an identity operation (e.g., $IXIX$ as opposed to $IXXI$).
+Our particular choice of logical operators is motivated by the fact that $\overline{XX}$ and $\overline{YY}$ (i.e., the logical operators we use to implement the LDD or NDD sequences) have a natural staggering of their physical $X$ and $Y$ gates.
+In other words, in the implementation of the logical operators comprising NDD, nearest-neighbor qubits are always interleaved with an identity operation (e.g., $IXIX$ as opposed to $IXXI$).
 This is critical for nearest-neighbor crosstalk cancellation [65].
@@
 The |$\ket{\Psi_{\pm}}$ Bell states are prepared similarly.
-Appendix B Creating robust logical sequences
-As discussed above, we use the logical operators $\overline{XX}=XIXI$ and $\overline{YY}=IYIY\equiv YIYI$ (along with identity operators) to generate the LDD groups ${\mathcal{G}}_{{\text{LXY4}}}=\{I,\overline{XX},\overline{YY},\overline{ZZ}\}$, and ${\mathcal{G}}_{{\text{LXX}}}=\{I,\overline{XX}\}$.
-Cycling over the group elements creates the four-pulse sequences LXY4 and LXX shown in the red boxes of Fig.
-1(d,e).
-Since the main source of logical errors is $ZZ$ crosstalk, we optimize LDD so that it cancels such errors.
-Thus, in the case of LXX, we insert a stabilizer every other pulse-step to consistently apply a staggered sequence to all qubits.
-These sequences correspond to applying logical $X$ and $Y$ operators to both logical qubits.
-To create the robust sequences, we mirror the original sequences (hence the eight-pulse sequences) and instead of $X$ and $Y$, we apply $\tilde{X}=R_{X}(-\pi)$ $\tilde{Y}=R_{Y}(-\pi)$, respectively, such that each sequence undergoes a robust pulse sequence at the physical level.
-We also apply staggered physical XY4 to each qubit as shown in Fig.
-1(f).
-This sequence is also inherently robust at the physical level since each qubit receives an XY4 sequence, which is robust to pulse errors [62].
-However, it is not a universal decoupling sequence at the logical level, as seen in the figure (i.e., it lacks the ability to decouple arbitrary single logical-qubit errors).
-Appendix C The encoding/unencoding Lemma
-Encoding of stabilizer codes is a familiar problem [91]. Here we present an explicit version of encoding circuits that appears to be simpler than other versions found in the literature [92, 93].
-C.1 Lemma and proof
-Lemma 1.
-For any $[[n,k,d]]$ stabilizer code it is always possible to construct a unencoding circuit $V^{\dagger}$ such that $V^{\dagger}$ converts any group of canonical logical operators and a corresponding stabilizer group into single-qubit Pauli $X$ and $Z$ operators acting just on the first $k$ physical data qubits and into single-qubit Pauli $Z$ stabilizer operators on the remaining $n-k$ ancilla qubits.
-Proof.
-Our proof is constructive and relies on the basic Clifford group identities $\text{CNOT}_{i,j}(X_{i}X_{j})\text{CNOT}_{i,j}=X_{i}$ and $\text{CNOT}_{i,j}(Z_{i}Z_{j})\text{CNOT}_{i,j}=Z_{j}$, where $\text{CNOT}_{i,j}$ denotes the controlled-NOT gate with qubit $i$ as the control and qubit $j$ as the target.
-Note that this identity allows us to reduce the number of $X$ operators.
-Consider first the case $k>0$.
-  1. 1.
-    Order the canonical logical operators and stabilizers as $\{\overline{X}_{1},\overline{Z}_{1},\dots,\overline{X}_{k},\overline{Z}_{k},S_{1},...,S_{n-k}\}$.
-  2. 2.
-    Ensure that $\overline{X}_{1}$ consists only of $X$ and $I$ physical operators: apply $H=(X+Z)/\sqrt{2}$ and $(X+Y)/\sqrt{2}$ to the qubits where $Z$ and $Y$, respectively, are present in the expression for $\overline{X}_{1}$.
-  3. 3.
-    Apply SWAP1,j if necessary to ensure $X_{1}$ is present in the expression for $\overline{X}_{1}$.
-  4. 4.
-    Apply $Z_{1}$ if necessary to cancel the sign if the sign was $-1$.
-  5. 5.
-    Apply CNOT1,j to eliminate all $X$ components of $\overline{X}_{1}$ except $X_{1}$.
-  6. 6.
-    Apply analogs of steps 2, 4, and 5 for $\overline{Z}_{1}$.
-  7. 7.
-    Now note that all the other operators contain $I_{1}$ (indeed, we have $\overline{X}_{1}=X_{1}$, $\overline{Z}_{1}=Z_{1}$, and all other operators commute with these two, hence they cannot contain $X_{1}$, $Z_{1}$, or $Y_{1}$), so the task is reduced to the case with $k:=k-1$ and $n:=n-1$; proceed recursively.
-Next, consider the case $k=0$. We have stabilizer operators $S_{1},...,S_{n}$ on $n$ physical qubits and wish to transform them to $Z_{1},...,Z_{n}$.
-  1. 1.
-    Apply analogs of previous steps 2-5 to $S_{1}$ to transform it into $Z_{1}$.
-  2. 2.
-    Recursively call this procedure to transform the remaining $n-1$ components of the remaining $n-1$ stabilizers to $Z_{2}$, …, $Z_{n}$.
-  3. 3.
-    Apply CNOT1,j gates to cancel the $Z_{1}$ components of the remaining $n-1$ stabilizers (if they had one).
-Note that they could not have had $X_{1}$ or $Y_{1}$ components because otherwise they would not commute with $S_{1}$.
-  4. 4.
-    Apply $X_{j}$ to cancel the sign of the remaining stabilizers if necessary.
-C.2 Illustration using the $[[4,2,2]]$ code
-As an example, consider the $[[4,2,2]]$ code.
-The stabilizer is $\mathcal{S}=\langle X^{\otimes 4},Z^{\otimes 4}\rangle$, and we choose the canonical logical operators as $\mathcal{L}=\langle\overline{X}_{1}=X_{1}X_{4},\overline{X}_{2}=X_{1}X_{2},\overline{Z}_{1}=Z_{1}Z_{2},\overline{Z}_{2}=Z_{1}Z_{4}\rangle$.
-The corresponding code basis states are
-|$\ket{\overline{00}}=\frac{1}{\sqrt{2}}\ket{0000}+\ket{1111}$ (9)
-|$\ket{\overline{10}}=\frac{1}{\sqrt{2}}\ket{0110}+\ket{1001}$
-|$\ket{\overline{01}}=\frac{1}{\sqrt{2}}\ket{0011}+\ket{1100}$
-|$\ket{\overline{11}}=\frac{1}{\sqrt{2}}\ket{1010}+\ket{0101}.$
-Following the steps in the proof of 1, we obtain the following series of transformations:
-$\overline{X}_{1}$ $=X_{1}X_{4}\stackrel{{\scriptstyle\text{CNOT}_{14}}}{{\longmapsto}}X_{1}{\longmapsto}X_{1}{\longmapsto}X_{1}{\longmapsto}X_{1}{\longmapsto}X_{1}$
-$\overline{Z}_{1}$ $=Z_{1}Z_{2}\longmapsto Z_{1}Z_{2}\stackrel{{\scriptstyle\text{CNOT}_{21}}}{{\longmapsto}}Z_{1}{\longmapsto}Z_{1}{\longmapsto}Z_{1}{\longmapsto}Z_{1}$
-$\overline{X}_{2}$ $=X_{1}X_{2}{\longmapsto}X_{1}X_{2}X_{4}{\longmapsto}X_{2}X_{4}\stackrel{{\scriptstyle\text{CNOT}_{24}}}{{\longmapsto}}X_{2}{\longmapsto}X_{2}{\longmapsto}X_{2}$
-$\overline{Z}_{2}$ $=Z_{1}Z_{4}\longmapsto Z_{4}{\longmapsto}Z_{4}{\longmapsto}Z_{2}Z_{4}\stackrel{{\scriptstyle\text{CNOT}_{42}}}{{\longmapsto}}Z_{2}{\longmapsto}Z_{2}$
-$S_{1}$ $=X^{\otimes 4}{\longmapsto}X_{1}X_{2}X_{3}{\longmapsto}X_{2}X_{3}{\longmapsto}X_{2}X_{3}X_{4}{\longmapsto}X_{3}X_{4}$
-$\quad\stackrel{{\scriptstyle H_{3}H_{4}}}{{\longmapsto}}Z_{3}Z_{4}\stackrel{{\scriptstyle\text{CNOT}_{34}}}{{\longmapsto}}Z_{3}$
-$S_{2}$ $=Z^{\otimes 4}{\longmapsto}Z_{2}Z_{3}Z_{4}{\longmapsto}Z_{2}Z_{3}Z_{4}{\longmapsto}Z_{3}Z_{4}{\longmapsto}Z_{3}Z_{4}$
-$\quad{\longmapsto}Z_{3}Z_{4}{\longmapsto}Z_{4}{\longmapsto}Z_{4}\ ,$ (10)
-where we have only indicated the transformations once, for the operator that induces the transformation per the proof of 1; e.g., $\text{CNOT}_{14}$ is induced by $\overline{X}_{1}$ but is applied to every operator in the first timestep, etc.
-Figure 8: Performance of LDD on for an encoded |$\ket{\Phi_{+}}$ Bell state on ibm_marrakesh for nine set of qubits.
-(a) shows the logical errors and (b) shows the postselection results.
-This device features tunable couplers that significantly reduce the crosstalk.
-Nevertheless, accumulation of dephasing errors over time (up to $\approx 55\leavevmode\nobreak\ \mu$s) lead to the appearance of logical errors.
-The same LDD sequence we used in the main text successfully suppresses these errors, as shown in (a).
-Additionally, LDD with postselection outperforms postselection alone, as shown in (b).
-Thus, the unencoding circuit, i.e., the circuit mapping every logical and stabilizer operator from the $[[4,2,2]]$ code to the corresponding operators of the trivial code, is
-$V^{\dagger}=\text{CNOT}_{34}H_{4}H_{3}\text{CNOT}_{24}\text{CNOT}_{42}\text{CNOT}_{21}\text{CNOT}_{14},$ (11)
-where $H_{i}$ is the Hadamard on the $i$th qubit.
-After unencoding, we obtain the trivial code with $\mathcal{S}_{0}=\langle Z_{3},Z_{4}\rangle$ and $\mathcal{L}_{0}=\langle X_{1},X_{2},Z_{1},Z_{2}\rangle$, that divides the code into $k=2$ data qubits and $n-k=2$ ancillas.
-In the trivial code, error correction is as follows: apply $IIXI$ when $IIZI=-1$ and $IIIX$ when $IIIZ=-1$.
-Applying $V^{\dagger}$ to these two recovery operators we find that to correct an undecoupled error in the original $[[4,2,2]]$ code, one should apply $IIIZ$ whenever $XXXX=-1$ and $XXXI$ (or $IIIX$, since they are related by a stabilizer element) when $ZZZZ=-1$.
+Appendix C Construction of our $[[4,2,2]]$ code subgroup-NDD sequence
+One of the limitations of our experimental setup is that physical $Z$ gates are unavailable: fixed-frequency transmons utilize virtual-$Z$ gates [96, 88].
+Such gates are problematic for DD [88].
+Here, now demonstrate that there are DD sequences satisfying the conditions of 1 which do not use any physical $Z$ pulses in the context of the $[[4,2,2]]$ code.
+To construct such a sequence, we first choose generators of $\mathcal{L}$ that do not involve physical $Z$’s.
+A simple example is $\mathcal{L}=\langle i,XIIX,IIXX,IIYY,YIIY\rangle$.
+One could follow the construction below with this choice.
+However, in order to improve the ability of the DD sequence to suppress $ZZ$ crosstalk, we make a different choice.
+First, multiply the original generators of $\mathcal{L}$ (that is, $XIIX,IIXX,IIZZ,ZIIZ$) by the stabilizers $XXXX$, $XXXX$, $XXXX$, and $ZZZZ$, respectively, to obtain $\mathcal{L}^{\prime}=\langle i,IXXI,XXII,XXYY,IZZI\rangle$.
+Then, after some group operations, we can write the same group using a different set of generators containing no $Z$’s: $\mathcal{L}^{\prime}=\langle i,XXII,IIYY,IYIY,XIXI\rangle$.
+The corresponding NDD sequence consisting of two repetitions of $XIXI$, $IYIY$, $XIXI$, $IIYY$, $XIXI$, $IYIY$, $XIXI$, and $XXII$, which implements the decoupling group /{±$\mathcal{L}^{\prime}/\{\pm 1,\pm i\}$, and involves no $Z$ gates.
+This sequence was obtained as follows: denote the generators of $\mathcal{L}$ (or $\mathcal{L}^{\prime}$) as $h_{1},h_{2},h_{3},h_{4}$ and choose any Gray code—a sequence of bitstrings $a_{0},\dots,a_{15}$ such that neighboring bitstrings differ in only a single digit.
+Then, pick $g_{j}=h^{a_{j}}$ (for $j=0,\dots,15$), where $a_{j}$ is interpreted as a multiindex (i.e., $g_{j}=\prod_{k=1}^{4}h_{k}^{(a_{j})_{k}}$).
+The pulses of the DD sequence are $g_{j}g_{j+1}$ for $j=0,\dots,15$.
 Appendix D Issues involving the implementation of logical state tomography using the $[[4,2,2]]$ code
@@
 For these reasons, we did not use logical state tomography in this work.
-Appendix E Logical dynamical decoupling with tunable couplers
+Appendix E Creating robust logical sequences
+As discussed in the main text, we use the logical operators $\overline{XX}=XIXI$ and $\overline{YY}=IYIY\equiv YIYI$ (along with identity operators) to generate the LDD groups ${\mathcal{G}}_{{\text{LXY4}}}=\{I,\overline{XX},\overline{YY},\overline{ZZ}\}$, and ${\mathcal{G}}_{{\text{LXX}}}=\{I,\overline{XX}\}$.
+Cycling over the group elements creates the basis for the four-pulse sequences shown in the red boxes of Fig.
+1(d,e).
+However, we add two additional components to enhance crosstalk cancellation and robustness.
+Since the main source of logical errors is $ZZ$ crosstalk, we modify LDD so that it cancels such errors.
+Thus, in the case of LXX, we insert a stabilizer every other pulse-step to consistently apply a staggered sequence to all qubits; in this manner, we move from LDD to NDD.
+These sequences correspond to applying logical $X$ and $Y$ operators to both logical qubits.
+To create the robust sequences, we mirror the original sequences (hence the eight-pulse sequences) and instead of $X$ and $Y$, we apply $\tilde{X}=R_{x}(-\pi)$, $\tilde{Y}=R_{y}(-\pi)$, respectively, such that each sequence undergoes a robust pulse sequence at the physical level.
+We also apply staggered physical XY4 to each qubit as shown in Fig.
+1(f).
+This sequence is also inherently robust at the physical level since each qubit receives an XY4 sequence, which is robust to pulse errors [62].
+However, it is not a universal decoupling sequence at the logical level, as seen in the figure (i.e., it lacks the ability to decouple arbitrary single logical-qubit errors).
+Figure 8: Performance of NDD on for an encoded |$\ket{\Phi_{+}}$ Bell state on ibm_marrakesh for a single set of qubits.
+(a) shows the logical errors and (b) shows the postselection results.
+This device features tunable couplers that significantly reduce the crosstalk.
+Nevertheless, accumulation of dephasing errors over time (up to $\approx 55\leavevmode\nobreak\ \mu$s) leads to the appearance of logical errors.
+The same NDD sequence we used in the main text successfully suppresses these errors, as shown in (a).
+Additionally, NDD with postselection outperforms postselection alone, as shown in (b).
+Appendix F Normalizer dynamical decoupling with tunable couplers
 Here we present additional experiments on ibm_marrakesh, which features tunable couplers [98] unlike the always-on $ZZ$ interaction in ibm_kyiv.
@@
 Nevertheless, a combination of dephasing and residual crosstalk accumulates, leading to logical errors.
-Figure 8 illustrates the performance of LDD sequences on this QPU, averaged over 9 different qubit sets.
-Fig.
-8(a) [equivalent to Fig.
+Figure 8 illustrates the performance of NDD sequences on this QPU, averaged over $9$ different qubit sets.
+Figure 8(a) [equivalent to Fig.
 3(b)] shows that without DD, logical errors remain unsuppressed.
-However, both logical DD sequences we employ successfully suppress these errors.
-Fig.
-8(b) [equivalent to Fig.
-5(a)] demonstrates that using only postselection, the fidelity averages to $\mathcal{F}\approx 56.23\%$, whereas combining postselection with LDD increases it to $\mathcal{F}\approx 91.73\%$.
-Optimizing LDD for specific types of tunable couplers to maximize the interplay between DD and couplers is an avenue for future work.
+However, both NDD sequences we employ successfully suppress these errors.
+Figure 8(b) [equivalent to Fig.
+5(a)] demonstrates that using only postselection, the fidelity averages to $\mathcal{F}\approx 56.23\%$, whereas combining postselection with NDD increases it to $\mathcal{F}\approx 91.73\%$.
+Optimizing NDD for specific types of tunable couplers to maximize the interplay between DD and couplers is an avenue for future work.